### v1

# Stratified Selective Sampling for Instruction Tuning with Dedicated Scoring Strategy

Abstract
Recent work shows that post-training datasets for LLMs can be substantially downsampled without noticeably deteriorating performance. However, data selection often incurs high computational costs or is limited to narrow domains. In this paper, we demonstrate that data selection can be both—efficient and universal—by using a multi-step pipeline in which we efficiently bin data points into groups, estimate quality using specialized models, and score difficulty with a robust, lightweight method. Task-based categorization allows us to control the composition of our final data—crucial for finetuning multi-purpose models. To guarantee diversity, we improve upon previous work using embedding models and a clustering algorithm. This integrated strategy enables high-performance fine-tuning with minimal overhead.
Stratified Selective Sampling for Instruction Tuning
with Dedicated Scoring Strategy
Paramita Mirza1, Lucas Weber1, Fabian Küch1 1Fraunhofer IIS {first.last}@iis.fraunhofer.de
1 Introduction
Data selection has long been a central challenge in machine learning, aimed at curating datasets that maximise model performance. This has gained particular significance in the era of large language models (LLMs) (Albalak et al., 2024). LLM training typically spans three stages—pretraining, instruction-tuning, and alignment—with data selection playing a crucial role in each (e.g. Grattafiori et al., 2024b). However, the training objectives differ between each phase, and the goals of data selection also vary accordingly. While in pretraining the selection mechanisms are relatively coarse, relying on heuristics or lightweight classifiers to filter out the most undesirable data (Rae et al., 2021; Lee et al., 2022; Weber et al., 2024; Penedo et al., 2024), the post-training phase of instruction-tuning and subsequent alignment requires, in comparison, more rigorous selection strategies to guide the model toward desirable behaviors (Longpre et al., 2023; Conover et al., 2023; Wang et al., 2023; Grattafiori et al., 2024b).
Recent research on instruction tuning (IT) underscores the critical importance of data quality over sheer data quantity, establishing the benefits of IT data selection beyond merely reducing costs (see e.g. Zhou et al., 2023a; Liu et al., 2025; Qin et al., 2025), as larger datasets not only yield diminishing returns as they expand, but they can have detrimental effects on performance, especially with synthetic data (Ge et al., 2024; Diddee and Ippolito, 2025). On the other hand, given the abundance of available instruction data, one has to make a selection decision.
Figure 1: An illustration of our three-stage pipeline: we first classify instructions into seven types, then score each sample for instruction difficulty ($f_{i}$) and response quality ($q_{i}$). We rank their combined scores ($p_{i}$) and sample top scoring examples within embedding-space clusters as well as overall, while maintaining fixed category proportions.
Manually curating effective instruction sets is both time-consuming and labour-intensive (Zhou et al., 2023a). A promising alternative is to automatically filter a small but diverse set of effective instructions from the extensive pool of available instruction data. Previous studies have proposed strategies to select “effective” instructions, focusing on three key dimensions of data assessment (Qin et al., 2025): i) difficulty/relevance, assessing whether a sample is trivial or challenging and how much it can therefore contribute to performance improvements, ii) quality, which refers to the usefulness and accuracy of responses, and iii) diversity, emphasising the scope of within-domain variability and cross-domain coverage of the data.
Prior work and its limitations.
Prior work highlights the value of complex or challenging examples for post-training (Li et al., 2024a; Liu et al., 2024a), but often relies on prompting LLMs to estimate complexity (Liu et al., 2024b; Lu et al., 2024) and lacks domain generalizability (Muennighoff et al., 2025). Similarly, quality scoring methods (e.g., Liu et al., 2024b; Ge et al., 2024) often ignore task-specific scoring needs (e.g., solutions’ correctness for math or coding problems), and some depend on large or proprietary models, raising cost and reproducibility issues (Chen et al., 2024). Domain-aware scoring strategies require a preprocessing step to bucket samples by task type. While some works mention this as part of data analysis (Ouyang et al., 2022a; Conover et al., 2023) or filtering pipelines (Grattafiori et al., 2024b), concrete methods and applications remain underexplored. Furthermore, as we detail in Section 2, only a few prior works consider all aspects of data assessment simultaneously (e.g., Liu et al., 2024b), and most evaluate on a single benchmark, limiting generalizability.
Our contributions.
In this paper, we propose a robust instruction-tuning data selection framework that simultaneously considers difficulty, quality and diversity, and we demonstrate its effectiveness across a wide range of benchmarks. Our key contributions are as follows: (i) a novel data sampling strategy for instruction-tuning that integrates task types (e.g., math, coding, generation), difficulty, and quality scores, while preserving both inter- and intra-class diversity through clustering and ranking; (ii) an efficient method for classifying IT data samples by task type; (iii) a new and efficient approach to estimate difficulty scores using model performance across diverse benchmarks; (iv) task-specific quality scoring strategies, including custom-designed scorers, particularly for constrained generation and coding tasks focusing on responses’ correctness; and (v) a large-scale evaluation spanning models of different families and sizes, with comparisons to strong baselines.
2 Related work
IT data selection.
Prior work on IT data selection generally falls into two categories based on how they assess sample difficulty, quality, and diversity: external scoring and model-inherent criteria. External scoring methods rely on: (i) handcrafted features such as coherence, grammaticality, naturalness and understandability (Cao et al., 2024); (ii) heuristics based on length or formatting (Zhao et al., 2024a; Muennighoff et al., 2025); and (iii) scores derived from LLMs of varying scales (LLM-scorers). Notable approaches leveraging LLM-scorers are: AlpaGasus (Chen et al., 2024), which prompts ChatGPT to score each data point; InsTag (Lu et al., 2024), which uses ChatGPT to generate open-ended tags for deriving complexity/diversity; Deita (Liu et al., 2024b), which fine-tunes LLaMA-1-13B on ChatGPT-annotated complexity/quality labels as deita-scorers; and CaR (Ge et al., 2024), which trains a 550M reward model to rank instruction-response pairs by quality. On the other hand, model-inherent criteria form another group of approaches that rely on signals from the target model itself, including: LESS (Xia et al., 2024), which estimates data influence using gradient information; SelectIT (Liu et al., 2024a), which measures uncertainty via token probability, prompt variation and multiple models’ assessments; SHED (He et al., 2024), which clusters data and estimates impact using Shapley values; instruction-following difficulty (IFD, Li et al., 2024a) measuring discrepancies between the model’s intrinsic generation capability and its desired response; and the approach by Li et al. (2024b), which evaluates sample utility based on how much it reduces loss when used as an in-context example.
Our approach combines both perspectives: we use domain-specific LLM-scorers to assess quality, and leverage various models’ performances across benchmarks to estimate difficulty—that is, how likely an average model is to fail on a given instruction. For diversity, our approach closely follows CaR (Ge et al., 2024), which clusters data points and samples one representative from each cluster.
IT data categorization.
Previous work has proposed various task types (e.g., open QA, brainstorming, creative writing), to guide IT data collection from human annotators (Ouyang et al., 2022a; Conover et al., 2023). Grattafiori et al. (2024b) finetuned Llama 3 8B for coarse-grained (e.g., mathematical reasoning) and fine-grained (e.g., geometry and trigonometry) topic classification to help filter low-quality samples, though they provide little detail on the methodology or intended downstream use. Other efforts impose tags or domain taxonomy on IT data to ensure diversity (Lu et al., 2024; Muennighoff et al., 2025). Dong et al. (2024) study how data composition across tasks affects model performance, but assume that ShareGPT only contains general alignment tasks—we found that nearly 30% of it actually consists of coding tasks. To our knowledge, no prior work leverages IT data categorization to apply task-specific scoring strategies during selection.
Utilities of IT data selection methods.
Several papers criticized the effectiveness and cost-efficiency of existing IT data selection strategies. Zhao et al. (2024a) show that simply selecting the longest responses can outperform more complex methods while being significantly cheaper and easier to implement. Similarly, Diddee and Ippolito (2025) find that many sophisticated methods barely outperform random sampling under realistic conditions, and emphasize the cost-performance trade-off. However, most of these comparisons are limited to a single-source sampling setup, whereas a more practical scenario involves selecting from a pool of IT data sources. Additionally, prior evaluations focus on LLaMA models with a few exceptions using Mistral Liu et al. (2025), leaving the generalizability of selection strategies across model families and sizes largely unexplored.
3 Methods
Our goal is to determine a subset of an arbitrarily large-scale instruction dataset that is effective when finetuning a given pre-trained base model. We define an effective dataset as one which achieves high performance on a given set of general evaluation tasks, while requiring relatively few model parameter updates. Formally, let $\mathcal{D}=\{d_{i}\}_{i=1}^{N}$ with $d_{i}=(x_{i},y_{i})$ be the full IT dataset of size $N$, where $x_{i}$ represents an input sequence (i.e. an instruction) and $y_{i}$ represents the corresponding output (i.e. desirable model response). We wish to select a subset $\mathcal{D}^{\prime}\subset\mathcal{D}$ of size $m$ (with $m\ll N$) that is most effective for instruction tuning.
Previous research has shown that i) instruction difficulty (see e.g. Li et al., 2024a; Liu et al., 2024a, b; Zhao et al., 2024b) ii) response quality (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b) and iii) diversity/composition (see e.g. Ge et al., 2024; Lu et al., 2024) are crucial for effective data selection for LLM finetuning. We address these three aspects in a three-step pipeline as illustrated in Figure 1:
  1. 1.
    Classification. We train a lightweight classifier $\pi_{c}$ to categorise all inputs $x_{i}$ in $\mathcal{D}$ into one out of seven categories, which we denote as $l\in\mathcal{L}$.
  2. 2.
    Scoring. Each input–output pair $(x_{i},y_{i})$ is scored using a category-specific quality scorer (yielding $f_{i}$) and a general-purpose difficulty scorer (yielding $q_{i}$); the results are combined into an overall preference score $p_{i}$.
  3. 3.
    Clustering + Ranking. Ultimately, we select the samples with the highest $p_{i}$ while using a clustering approach to maintain diversity and minimise redundancies in $\mathcal{D}^{\prime}$.
We detail each step in the remainder of this section.
3.1 Instruction Classification
Inspired by the use-case categories defined in Ouyang et al. (2022b), we establish the following task categories $\mathcal{L}$ for samples in instruction tuning datasets: i) Math, from simple calculation to problems requiring multi-step reasoning; ii) Coding, code generation tasks or programming-related question answering; iii) Generation, textual generation tasks including roleplaying, summarizing and rewriting passages; iv) Reasoning, questions requiring deductive/logical reasoning; v) Brainstorming, information-seeking and recommendation questions that require inductive reasoning, including classification tasks; vi) Factual QA, factual questions with simple facts as answers; and vii) Extraction, tasks requiring structured/answer extraction from textual contexts. We let two human annotators classify 80 samples from the popular MT-Bench dataset (Zheng et al., 2023a) and achieve high inter-annotator agreement (Cohen’s Kappa = 0.8635), demonstrating the discriminability of our categories. We compare two different approaches to build a classification model: i) LLM annotator and ii) SetFit classifier.
With the LLM-annotator approach (Wei et al., 2022), we prompt instruction categorization by listing categories with brief explanations, followed by “What is the category of the following task?” (see Figure 12, Appendix A.1). Meanwhile, SetFit (Tunstall et al., 2022) is a few-shot learning method that tunes Sentence Transformers (Reimers and Gurevych, 2019) on labelled input pairs in a contrastive, Siamese manner. We manually identify approximately 250 samples strongly associated with each category (see Table 3, Appendix A.1) to train the SetFit classifier; hyperparameters are detailed in Appendix A.1.
For evaluation, we use the manually annotated MT-Bench dataset as the test set, assessing accuracy, macro F1-score, and Cohen’s Kappa agreement with human judgment (see Table 1). While zero-shot prompting with GPT-4o performs best, we choose the SetFit classifier for our pipeline due to its comparable performance and higher efficiency than larger LLMs like GPT-4o and Falcon3-10B-Instruct (Team, 2024b).
Approach LLM annotator/embedding model Acc F1 Cohen’s Kappa Zero-shot GPT-4o 0.88 0.86 0.85 prompting tiiuae/Falcon3-10B-Instruct 0.85 0.84 0.82 meta-llama/Llama-3.1-8B-Instruct 0.76 0.65 0.71 SetFit NovaSearch/stella_en_400M_v5 $\dagger$ 0.85 0.81 0.82 classifier Lajavaness/bilingual-embedding-large 0.82 0.78 0.79 NovaSearch/stella_en_1.5B_v5 0.66 0.60 0.60
Table 1: Evaluation results on instruction categorization. $\dagger$ denotes the chosen approach and model for our data selection pipeline.
3.2 Scoring
Domain-agnostic difficulty scorer.
Previous research suggests that difficulty (sometimes referred to as complexity) matters for data selection (see e.g. Liu et al., 2024b; Cao et al., 2024; Zhao et al., 2024b; Muennighoff et al., 2025), with more challenging data generally resulting in better model performance. However, existing difficulty metrics either often lack generality across domains (e.g. length of reasoning trace in response in Muennighoff et al., 2025) or are strongly influenced by spurious features (e.g. the widely used deita-complexity is strongly biased towards long sequences; see Figure 7 in Appendix A.2.1; Liu et al., 2024b).
Our goal is to train a general and robust difficulty scorer for our data selection pipeline that predicts how likely it is for an average model to solve a data point incorrectly, independent of its category $l$.
To source the training set $\mathcal{D}^{\text{diff}}$ for such a scorer, we collect 20k instruction-response pairs, evenly distributed across categories $l_{i}\in\mathcal{L}$ (data sources are listed in Table 4; proportions are shown in Figure 8). We evaluate every item with a heterogeneous pool of 18 instruction-tuned LLMs (see Table 5). Subsequently, we apply multiple preprocessing steps to the model scores: First, we normalise the item scores to the interval $[0,1]$ and remove items with a score of 0 across all models, as they are likely to contain noise or annotation errors. To mitigate potential skews in the model performance distribution, we convert the absolute scores into relative deviations with respect to the model’s mean performance on the corresponding category source dataset, by subtracting the average from the absolute score on each item. Ultimately, a difficulty target for every item is obtained by averaging over the model pool. We fine-tune a Qwen-3-8B backbone (Team, 2025) with a single-layer regression head to minimise the mean-squared error on this dataset (training details can be found in Appendix A.2.3).
Response quality scorer.
Prior work shows that sample selection strategies based on response length or quality, as judged by external models, lead to better instruction tuning datasets (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b). However, these strategies underestimate the diverse problem types within instruction tuning dataset, which may require different evaluation criteria. For example, the quality of responses to math and coding problems is heavily dependent on solution correctness, while constrained generation requires evaluation of adhered constraints. In this work, we designate a dedicated quality scorer for each category defined in Section 3.1, focusing on i) mathematical reasoning traces for Math ($q_{\mathrm{math}}$), ii) code snippets for Coding ($q_{\mathrm{code}}$), and iii) instruction-following capability for Generation and Brainstorming ($q_{\mathrm{if}}$). For Reasoning, Factual QA and Extraction, we employ the deita-quality scorer [^1] https://huggingface.co/hkust-nlp/deita-quality-scorer ($q_{\mathrm{deita}}$, Liu et al., 2024b), a finetuned LLaMA-13B for quality assessment,
Process reward model.
Process reward models (PRMs, Lightman et al., 2023; Uesato et al., 2022) are trained to verify steps in reasoning traces as they are common in mathematical reasoning. We score Math data points using Qwen2.5-Math-PRM-7B (Zhang et al., 2025b). We find double linebreaks and—if no double linebreaks present—single linebreaks as a delimiter to be a good heuristic to separate reasoning steps. As a reasoning trace breaks with a single erroneous step, we aggregate scores by taking the minimal score out of all steps within each trace, as the $q_{\mathrm{math}}$ score.
Code quality scorer.
We design a quality scoring framework for Coding samples, drawing inspiration from Wadhwa et al. (2024). For each data point $(x_{i},y_{i})$, we leverage code-oriented LLMs to: (i) assess the functional correctness of the code snippet in $y_{i}$ with respect to the problem $x_{i}$, and (ii) produce a revised version that improves or fixes the original code (see Figure 13, Appendix A.3). The resulting score, $q_{\mathrm{code}}$, is based on the normalized Levenshtein similarity between lines of the original ($lo_{0},...,lo_{n}$) and revised ($lr_{0},...,lr_{m}$) code: $nls=(\max(n,m)-\mathrm{lev}(lo,lr))/\max(n,m)$, where $\mathrm{lev}(lo,lr)$ is the line-level Levenshtein distance. If the original code is functionally correct, we set $q_{\mathrm{code}}=nls$; otherwise, $q_{\mathrm{code}}=nls/2$. If no code snippet is present, we assign $q_{\mathrm{code}}=0.5$ if $y_{i}$ is judged correct, and 0.0 otherwise.
To evaluate this scoring method, we use a 1K-sample test set from LiveCodeBench (Jain et al., 2024a), containing coding problems and LLM-generated responses. [^2] https://huggingface.co/spaces/livecodebench/code_generation_samples Using Qwen/Qwen2.5-Coder-14B-Instruct as the reviewing model, our framework achieves 70% accuracy and a 0.412 Pearson correlation with binary correctness labels (see Appendix A.3 for details).
Instruction-following scorer.
Taking inspiration from the IFEval benchmark (Zhou et al., 2023b) which defines “verifiable constraints” such as length (“400 or more words”) and keyword (“without using the word sleep”) constraints, we design a response quality scorer based on the fraction of expressed constraints ($\mathcal{C}_{exp}$) adhered by the response ($\mathcal{C}_{true}$). First, we use an LLM annotator to identify $\mathcal{C}_{exp}$, which comprises (span, constraint type) pairs $\{(s_{i},c_{i})\}^{n_{exp}}_{i=1}$, with $s_{i}$ represents the textual span found and $c_{i}$ is the corresponding constraint label. For example, given the prompt “Write a funny blog post with 400 or more words about the benefits of sleeping in a hammock, without using the word sleep.”, $\mathcal{C}_{exp}=$(400 or more words, length), (without using the word “sleep”, keyword avoided), (funny blog post, writing type)}. A list of considered constraint types is shown in Figure 14 (Appendix A.4). Next, $\mathcal{C}_{exp}$ is passed to a constraint checker module, which performs two steps:
  1. 1.
    Heuristic verification: We verify length, letter case, punctuation and keyword constraints, by adapting the IFEval verification script.
  2. 2.
    LLM-judge verification: We ask an LLM judge to assess constraints that cannot be verified heuristically (e.g., “Does the following text follow the [writing type] constraint of [funny blog post]?”).
This yields $\mathcal{C}_{true}=\{(s_{j},c_{j})\}^{n_{true}}_{j=1}$, with which we compute the quality score as $q_{\mathrm{if}}=n_{true}*(n_{true}/n_{exp})$, giving more incentives to responses adhering to more constraints. If $\mathcal{C}_{exp}$ is empty, we ask an LLM judge to evaluate whether the response i) addresses the user’s intent, while ii) respecting any constraints expressed in the prompt, and to provide a final $score$ (1–10), which we use to compute the score as $q_{\mathrm{if}}=score/10$.
Our analysis with the IFEval benchmark dataset containing sample responses from ten models as our test bed (see Table 8, Appendix A.4), shows that Qwen3-14B (Team, 2025) outperformed other medium-sized Instruct-LLMs as both LLM annotator and judge (see Table 9, Appendix A.4). It achieved a macro F1-score of 0.86 for identifying expressed constraints, a Pearson correlation coefficient of 0.523 at the instance-level, and 0.995 at the model-level, where it effectively replicated the IFEval model ranking.
Overall preference scores.
For each $(x_{i},y_{i})\in\mathcal{D}$, we compute the preference score $p_{i}=f_{i}\cdot q_{i}$, where $f_{i}$ and $q_{i}$ are the difficulty and quality scores, respectively. Each of them is normalized using min-max scaling, with the 1st and 99th percentiles as the minimum and maximum values across all samples in $\mathcal{D}$. The quality scores are normalized per scorer ($q_{math}$, $q_{code}$, $q_{if}$ and $q_{deita}$) as they have differing ranges. For multi-turn conversations, where each data point $d_{i}$ consists of a sequence of turns $\{(x_{0},y_{0}),...,(x_{T},y_{T})\}$, we assign a category $l_{t}$ to each turn and determine the conversation-level category $l_{i}$ by either selecting the most frequent category or defaulting to the first one. We then compute turn-level difficulty and quality scores, $f_{t}$ and $q_{t}$, based on their respective categories $l_{t}$. These are averaged across all turns to yield the overall conversation-level scores $f_{i}$ and $q_{i}$. [^3] For a given conversation, quality scores are averaged separately for each category; $q_{i}$ is then selected according to the main category $l_{i}$.
3.3 Sampling
Greedily choosing the highest-scoring samples often leads to redundancy in some domains and under-representation in others. Thus, maintaining diversity in the final dataset $\mathcal{D}^{\prime}$ is essential. Since diversity is a property of the dataset as a whole—not of individual samples—selection should consider the dataset globally (e.g., Ge et al., 2024), rather than relying on iterative, sample-by-sample strategies (e.g., Liu et al., 2024b; Bukharin et al., 2024). Diversity can be promoted top-down by balancing category proportions (see e.g. Grattafiori et al., 2024b; Dong et al., 2024), or bottom-up by ensuring sufficient semantic dissimilarity among selected samples (e.g., Liu et al., 2024b; Ge et al., 2024; Lu et al., 2024).
[…]
4 Experiments
4.1 Experimental Setup
Datasets.
We use the IT datasets detailed in Table 2. In our experiments, the full IT dataset $D$ is the aggregation of all listed datasets.
Dataset #samples (#turns) HuggingFaceH4/ifeval-like-data 5K vicgalle/alpaca-gpt4 52K nvidia/OpenMathInstruct-2555Without augmented problems. 52K ai2-adapt-dev/flan_v2_converted 90K openbmb/UltraInteract_sft (Coding) 115K WizardLMTeam/WizardLM_evol_instruct_V2_196K 143K theblackcat102/sharegpt-english 50K (392K) microsoft/orca-agentinstruct-1M-v1666Randomly sampled. 200K (903K) all 707K (1.75M)
Table 2: Instruction tuning dataset overview.
Figure 2: Category proportions of different sampling strategies for 100k samples.
Models.
While most experiment results use finetuned Mistral-7B-v0.3, we also demonstrate generalization across models of varying sizes and families: i) tiiuae/Falcon3-10B-Base, ii) meta-llama/Llama-3.1-8B, iii) mistralai/Mistral-7B-v0.3, iv) Qwen/Qwen2.5-3B and v) HuggingFaceTB/SmolLM2-1.7B.
(a) Aggregation models & benchmarks
(b) Aggregation benchmarks
(c) Aggregation models
Figure 3: Performance gains over the base model for different sampling strategies with 100k samples. We aggregate the results (a) across all 5 tested models and all 13 benchmarks; (b) across all benchmarks separated by models; (c) across all models separated by benchmarks.
Finetuning.
We fine-tune each base model using the SFTTrainer from the Transformer Reinforcement Learning (TRL) library [^7] https://huggingface.co/docs/trl/sft_trainer (von Werra et al., 2020). Model-specific hyperparameters (e.g., learning rate) are selected for each model family and detailed in Appendix A.6. We also employ NEFTune (Jain et al., 2024b), a technique that improves the performance of chat models by injecting noise into embedding vectors during training.
Baselines.
We compare against the following baseline methods for constructing $D^{\prime}\subset D$: (i) Random, where we sample uniformly at random, (ii) Longest (Zhao et al., 2024a), where we include samples having the longest responses and (iii) Deita (Liu et al., 2024b), which ranks data points according to complexity and quality scores ($o_{deita}$ and $q_{deita}$, resp.), and iteratively builds $D^{\prime}$ by adding samples that are dissimilar to those already selected.
Evaluation.
We evaluate the success of our instruction tuning experiments by testing Models on a suite of benchmarks, covering a broad spectrum of model capabilities and giving a holistic picture of the tuning success. All tested benchmarks are listed in Table 11 in Appendix A.7.
4.2 Main Results
For every selection strategy we construct a subset $\mathcal{D}^{\prime}={d}_{i=1}^{m}\subset\mathcal{D}$ with size $m=100,000$ and subsequently fine-tune each of the five base models on $\mathcal{D}^{\prime}$. In addition to the baseline strategies described in Section 4.1, we investigate four principled variants that exploit the pipeline components introduced in Section 3.2:
  1. i)
    Quality—sort $\mathcal{D}$ by $q_{i}$ and select the top $m$ items;
  2. ii)
    Difficulty—sort by $f_{i}$ and select the top $m$ items;
  3. iii)
    Combination—sort by the preference score $p_{i}=f_{i}\cdot q_{i}$ and select the top $m$ items;
  4. iv)
    Combination++—as in combination, but with sampling via the clustering-and-quota procedure from Section 3.3.
Figure 2 shows task-type distributions across sampling strategies, including baselines. Figure 3(a) reports average performance gains over the untuned base model for all evaluation benchmarks. The full pipeline (combination++) achieves the highest overall performance, surpassing every baseline and even the model trained on the entire source set with $|\mathcal{D}|>707\text{k}$.
To examine robustness, we break the results down by model size (Figure 3(b)). Combination++ yields the most consistent gains of all tested conditions for all five bases and is outperformed by full-data training in only two cases, confirming that the proposed sampling approach transfers well across models. A benchmark-wise analysis (Figure 3(c)) shows similarly stable improvements, with combination++ delivering consistent performance throughout.
(a) MT-Bench
(b) Alpaca Eval 2.0
Figure 4: Performance on LLM-as-a-judge benchmarks.
For completeness, we also evaluate our approach on benchmarks using LLM-as-a-judge. Due to the high cost of the proprietary LLM-judges, we limit ourselves to evaluating models trained on Mistral-7B-Base under a single experimental condition. As shown in Figure 4, for $m=100,000$, combination++ performs best in the length-controlled AlpacaEval 2.0 (Dubois et al., 2024), and matches random sampling on MT-bench (Zheng et al., 2023b).
Figure 5: Average benchmark score gains over the base model, when finetuning Mistral-7B-Base with different dataset sizes sampling using various strategies.
Robust downscaling
We next test whether the full pipeline’s benefits persist when the target size $m$ is varied. Fixing the base model to Mistral-7B-Base, we sample additional datasets $\mathcal{D}^{\prime}$ of 1k, 10k, 25k, 50k items with combination++ and all baselines. Figure 5 demonstrates that our method outperforms the alternatives across every scale: notably, with only $m=50,000$ (7% of the source data), it surpasses the performance resulting from full-data training. For benchmark-specific plots, see Appendix A.8.1.
Figure 6: (a) Category proportions of $\mathcal{D}_{\text{skewed}}$ and the 25k subsets sampled with various selection strategies; (b) Average benchmark scores, when finetuning Mistral-7B-Base with data sampled from the skewed distribution using various strategies.
We, again, evaluate on LLM-as-a-judge benchmarks under limited setups. As before, combination++ performs comparably to other strategies on MT-bench, while outperforming them on AlpacaEval for $m=5,000$ and on average performance for $m=25,000$.
The surprising insensitivtiy of MT-Bench to sampling strategies may stem from the consistent presence of GPT-4-generated responses in every sampled datasets (see Figure 10, Appendix A.5), which is known to strongly influence MT-Bench outcomes (Panickssery et al., 2024; Stureborg et al., 2024; Wataoka et al., 2024).
Robustness to skewed data distributions
We now probe how selection strategies perform when the candidate pool $\mathcal{D}$ is itself highly imbalanced. To simulate this, we create a skewed source set $\mathcal{D}_{\text{skewed}}\subset\mathcal{D}$ by randomly selecting two categories $l_{1},l_{2}\in\mathcal{L}$ and retaining only items with labels $l_{i}\in{l_{1},l_{2}}$, along with a small (4%) residue from all other categories. This yields a strongly biased data distribution of size $N_{\text{skewed}}=366\text{k}$ (see Figure 6(a) bottom, for the resulting category distribution).
From $\mathcal{D}_{\text{skewed}}$, we sample $m=25\text{k}$ items with the same strategies as previously and fine-tune Mistral-7B-Base on each $\mathcal{D}_{\text{skewed}}^{\prime}$. We expected diversity-aware sampling strategies like Deita and combination++ to mitigate bias by enforcing either semantic spread and explicit quotas. However, none of the strategies show a clear advantage, with random sampling performing with a slight, insignificant edge. On closer inspection of the category distributions of the sampled data in Figure 6(a), we find that sampling strategies without explicitly encouraging diversity, such as quality and difficulty, mostly maintain the bias of $\mathcal{D}_{\text{skewed}}$, while Deita exacerbates it. Combination++ successfully debiases the sample; however, this balancing does not translate into improved model performance.
5 Discussion and Conclusion
Effective data selection for instruction tuning is increasingly important as we handle the increasing amount of available data of mixed quality. Nevertheless, the results of data selection pipelines have often been brittle, struggling to generalise across training setups (Diddee and Ippolito, 2025; Zhao et al., 2024a).
In this paper, we address this challenge by explicitly covering the whole spectrum of possible instruction domains and tailoring scoring strategies that are apt to judge the utility of samples in those domains. We show the robustness of our approach by testing it across diverse settings (model families, scales, and sample sizes) and evaluating it on a wide range of common benchmarks. Our experiments provide further evidence for the necessity of data selection, as our sampling not only significantly reduces the amount of data required for training, but also outperforms setups trained on the full source data. Despite its more elaborate design—including multiple scoring methods—our pipeline remains computationally efficient due to targeted scoring.
While our pipeline outperforms baselines in almost all settings, it shows no significant gains when applied to a strongly skewed source distribution. We caution against over-interpreting this outcome, as we only evaluate on a single type of distributional bias (toward two of seven categories). A more complete robustness assessment would require testing across various bias configurations, which is computationally infeasible and thus left for future work.
Limitations
Difficulty scorer.
A major issue in collecting data for our difficulty scorer is a certain unreliability in the evaluation of model responses. Previous research shows that evaluations oftentimes have weak robustness to e.g. the prompt formatting (Sclar et al., 2024; Weber et al., 2023a, b; Polo et al., 2024), bias of LLMs-as-a-judge (Panickssery et al., 2024; Stureborg et al., 2024; Wataoka et al., 2024) or errors in post-processing (such as issues in extracting the answer from the model response). For example, GPT4o showed overall weaker performance on some of our evaluated subsets than some small open-source models. Upon closer inspection, we encountered that – while providing the correct answer – GPT4o generally tends not to follow the formatting of the given few-shot examples and rather responds in an open-form manner, resulting in failing the tight response search masks of the used evaluation frameworks. While we try to mitigate this issue as much as possible, we cannot guarantee that difficulty scores exactly reflect a model’s capacity to solve a given data point.
Quality scorer.
We did not develop nor have dedicated quality scorers for samples belonging to Reasoning, Factual QA and Extraction categories. However, we hypothesis that process reward models (PRMs) could be adapted to Reasoning tasks beyond math, such as spatial (e.g., Wu et al., 2024) and deductive reasoning (e.g., Seals and Shalin, 2024), given appropriate training data. Factuality assessment is a long-standing research problem that has become more relevant in the era of LLMs. Wei et al. (2024) introduced SAFE (Search-Augmented Factuality Evaluator), an LLM-agent-based method for automatically assessing long-form factuality in model response. Although significantly cheaper than human annotators (up to 20×20$\times$20 ×), SAFE still incurs costs of \$20–\$40 per 100 prompt-response pairs. For Extraction tasks, future work might draw inspiration from recent advances in RAG evaluation (Yu et al., 2025). These directions, however, are beyond the scope of this paper.
Robustness to skewed data.
Our results indicate that we cannot get performance gain in setups with very skewed data. However, the experimental setup is very limited, testing only skew in a single variation and more experiments are needed to draw conclusions. Further, there are many other skews that can occur in practice that we did not evaluate our sampling pipeline against.
Appendix A Appendix
A.1 Instruction Categorization – details
Category Training data (subset) #Samples Math nvidia/OpenMathInstruct-2, AI-MO/NuminaMath-CoT 250 Coding openbmb/UltraInteract_sft (Coding), microsoft/orca-agentinstruct-1M-v1 (code), 253 HuggingFaceH4/no_robots (Coding), lissadesu/codeqa_v3 Generation HuggingFaceH4/no_robots (Generation, Rewrite, Summarize), HuggingFaceH4/ifeval-like-data, 250 declare-lab/InstructEvalImpact (Creative, Professional), iamketan25/roleplay-instructions-dataset Extraction HuggingFaceH4/no_robots (Closed QA, Extract) 250 Factual QA HuggingFaceH4/no_robots (Open QA), basicv8vc/SimpleQA 255 Brainstorming declare-lab/InstructEvalImpact (Informative, Argumentative), HuggingFaceH4/no_robots (Brainstorming, Classify), 250 matt-seb-ho/WikiWhy Reasoning renma/ProofWriter, hitachi-nlp/ruletaker, lucasmccabe/logiqa, lucasmccabe/logiqa, tasksource/strategy-qa 250
Table 3: Training data overview for SetFit classifier.
We present the zero-shot prompt for classifying instructions in Figure 12. As for training the SetFit classifier, we leverage the training data detailed in Table 3, and we employ the following hyperparameters. Note that training with SetFit consists of two phases behind the scenes: finetuning embeddings and training a differentiable classification head. As a result, some of the training arguments can be tuples, where the two values are used for each of the two phases, respectively.
  • batch_size=(16, 2)
  • num_epochs=(1, 15)
  • end_to_end=True (train the entire model end-to-end during the classifier training phase)
  • body_learning_rate=(2e-5, 1e-5), the second value is the learning rate of the Sentence Transformer body during the classifier training phase
  • head_learning_rate=1e-4
  • max_steps=500
A.2 Difficulty Scorer – details
A.2.1 Correlations of difficulty/complexity metrics with input length
Figure 7: Relation between complexity/difficulty scores and length of the input sequence. The length of the input sequence should, in most cases, be unrelated to the difficulty of resolving it. Any correlation is therefore spurious.
As the length of an instruction might in some cases reflect the difficulty of the datapoint, there is no causal relationship between the two. A scorer that is trained to predict difficulty should therefore not rely on input length as a feature to rely its prediction upon. We show in Figure 7 how deita complexity is strongly correlated with input length, while the difficulty scorer we present in this paper is not. While this is not conclusive evidence that the complexity scorer is relying on input length as a feature in its prediction, these results are an indicator that it might. Further investigation is required to determine whether this relationship is causal.
A.2.2 Data generation difficulty scorer – details
Figure 8 show the proportions of different instruction-response pairs that we use a training data for our difficulty scorer. The subsequent Table 4 shows the respective data source from which we obtained the data, as well as the type of evaluation metric that we use to evaluate the 18 LLMs from Table 5.
Figure 8: Setfit proportions of training data for difficulty scorer
Dataset Subset Split $n$ samples Eval. metric
GSM8K (Cobbe et al., 2021) Default train 1192 exact match
Math (Hendrycks et al., 2021c) Algebra train 322 exact match
Counting & Probability train 264 exact match
Geometry train 220 exact match
Intermediate Algebra train 233 exact match
Number Theory train 314 exact match
Prealgebra train 348 exact match
Precalculus train 238 exact match
IFEval-like Default - 1990 instance level loose acc
MBPP (Austin et al., 2021) Default train & val 448 pass@1
OpenBookQA (Mihaylov et al., 2018) Default train 299 acc
ARC (Clark et al., 2018) Challenge train 464 acc
bAbI (Dodge et al., 2016) Default train 224 exact match
CommonsenseQA (Talmor et al., 2019) Default train 248 acc
CoQA (Reddy et al., 2019) Default train 380 F1
DROP (Dua et al., 2019) Default train 383 F1
FLD (Morishita et al., 2023) Default train 739 exact match
Logical Formula Default train 750 exact match
HeadQA (Vilares and Gómez-Rodríguez, 2019) English train 689 acc
Spanish train 703 acc
JSONSchemaBench (Geng et al., 2025) Easy train 200 schema compliance & json validity
Medium train 198 schema compliance & json validity
Hard train 179 schema compliance & json validity
LogiQA (Liu et al., 2020) LogiEval train 490 exact match
LogiQA2 train 416 acc
MLQA (Lewis et al., 2019) 49 lang. combinations val 715 F1
TriviaQA (Joshi et al., 2017) Default train 1389 exact match
OpenAssistent (Köpf et al., 2023) Default train 5318 LLM-as-a-judge
APPS (Hendrycks et al., 2021a) introductory train 422 pass@1
interview train 303 pass@1
competition train 289 pass@1
CONALA (Yin et al., 2018) Default train 97 Bleu
Table 4: Datasets used in difficulty scorer training.
Model Type Size
SmolLM2-135M-Instruct (Allal et al., 2025) Univ. 135M
SmolLM2-360M-Instruct (Allal et al., 2025) Univ. 360M
Qwen2.5-0.5B-Instruct (Team, 2024a) Univ. 0.5B
Qwen2.5-Math-1.5B-Instruct (Yang et al., 2024) Math 1.5B
Qwen2.5-Coder-1.5B-Instruct (Hui et al., 2024) Code 1.5B
Qwen2.5-1.5B-Instruct (Team, 2024a) Univ. 1.5B
SmolLM2-1.7B-Instruct (Allal et al., 2025) Univ. 1.7B
Qwen2.5-Math-7B-Instruct (Yang et al., 2024) Math 7B
Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) Code 7B
Qwen2.5-7B-Instruct (Team, 2024a) Univ. 7B
Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) Univ. 7B
Qwen2.5-14B-Instruct (Team, 2024a) Univ. 14B
Mistral-Small-24B-Instruct-2501 (Mistral AI, 2025) Univ. 24B
Qwen2.5-32B-Instruct (Team, 2024a) Univ. 32B
Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) Code 32B
Qwen3-32B (Team, 2025) Univ. 32B
gpt-4o-mini-2024-07-18 (Hurst et al., 2024) Univ. unk
gpt-4o-2024-08-06 (Hurst et al., 2024) Univ. unk
Table 5: Models that were evaluated to obtain difficulty scores for our difficulty scorer training set.
During the data collection for the difficulty scorer, we collect training sets from different benchmarks as well as data points from OpenAssistent (Köpf et al., 2023) for examples of open-ended generation. We evaluate these samples using the existing evaluation frameworks LM-evaluation harness (Gao et al., 2024), big-code evaluation harness (Ben Allal et al., 2022) and FastChat (Zheng et al., 2023b). For the LLM-as-a-judge evaluation, we use GPT4o as a judge and evaluate only a subset of 4 LLMs from Table 5 that we deem representative in terms of capabilities (gpt-4o-2024-08-06, Mistral-Small-24B-Instruct-2501, Mistral-7B-Instruct-v0.3 and SmolLM2-1.7B-Instruct).
A.2.3 Training the difficulty scorer – details
We equip regular CausalLLMs with a regression head by pooling the final dense layers and adding a linear projection to a scalar output, and finetune these models on the difficulty scoring problem. During finetuning, we use the hyperparameters detailed in Table 6. We finetune four different base models as difficulty scorer (Llama-3.1-8B (Grattafiori et al., 2024a), Qwen-2.5-7B (Team, 2024a), Qwen3-4B and Qwen3-8B (Team, 2025)) and converge on Qwen3-8B as it produces the best performance on our i.i.d. evaluation data.
Hyperparameter Value
batch_size 1
gradient_accumulation 16
learning_rate 1e-5
lr_scheduler_type linear
num_train_epochs 8
warmup_steps 100
max_seq_length 2048
weight_decay 0.01
neftune_noise_alpha 10
Table 6: Hyperparameter details for training the difficulty scorer
A.3 Code-quality Scorer – details
The prompt used by LLM-annotator/judge for deriving code-quality scores can be found in Figure 13. Table 7 presents the evaluation results of various LLMs as the code reviewer.
LLM reviewer acc Pearson’s $r$ deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct 0.64 0.278 deepseek-ai/DeepSeek-R1-Distill-Qwen-14B 0.66 0.389 Qwen/Qwen2.5-Coder-14B-Instruct $\dagger$ 0.70 0.412 Qwen/Qwen3-14B (non-coding LLM) 0.70 0.411
Table 7: Evaluation results of various LLMs as code reviewer (acc), and Pearson correlation coefficient ($r$) of resulting code quality scores with LiveCodeBench benchmarks binary correctness labels. $\dagger$ denotes the chosen LLM annotator/judge for our instruction-following scorer.
A.4 Instruction-following Scorer – details
Prompts used by LLM-annotator/judge for deriving instruction-following scores can be found in Figure 14, 15 and 16. As our test dataset, we collected responses from 10 LLMs on the IFEval benchmark (see Table 8), available on open-llm-leaderboard’s dataset collection of evaluation details. [^8] https://huggingface.co/open-llm-leaderboard Table 9 presents the evaluation results of various LLMs as the annotator/judge.
Model IFEval score
meta-llama/Llama-3.3-70B-Instruct 0.90
Qwen/Qwen2.5-14B-Instruct-1M 0.84
allenai/Llama-3.1-Tulu-3-70B-SFT 0.81
tiiuae/Falcon3-7B-Instruct 0.76
ibm-granite/granite-3.1-8b-instruct 0.72
microsoft/Phi-3-medium-128k-instruct 0.60
abacusai/Smaug-34B-v0.1 0.50
Qwen/Qwen2.5-32B 0.41
google/gemma-1.1-2b-it 0.31
databricks/dolly-v2-7b 0.20
Table 8: Performance of 10 considered LLMs on the IFEval benchmark.
LLM annotator/judge macro Pearson’s $r$ F1 instance-level model-level Qwen/Qwen2.5-7B-Instruct 0.80 0.503 0.986 meta-llama/Llama-3.1-8B-Instruct 0.83 0.454 0.982 tiiuae/Falcon3-10B-Instruct 0.84 0.515 0.969 Qwen/Qwen3-14B $\dagger$ 0.86 0.523 0.995
Table 9: Evaluation results of various LLMs as annotator/judge on identifying expressed constraints (macro-F1), and Pearson correlations coefficient ($r$) of resulting instruction-following scores with IFEval benchmarks scores at both instance-level and model-level. $\dagger$ denotes the chosen LLM annotator/judge for our instruction-following scorer.
Figure 9: Category proportions of IT datasets used in the experiments.
Figure 10: Source dataset proportions of different sampling strategies for 100k samples.
A.5 Datasets – details
Figure 9 shows SetFit classification results for IT datasets used in the experiments. Whereas Figure 10 shows the composition of source datasets for various sampling strategies. Deita is basically dominated by only two datasets: microsoft/orca-agentinstruct-1M-v1 (200k sampled) and WizardLMTeam/WizardLM_evol_instruct_V2_196K.
A.6 Finetuning – details
We finetune all models in our experiments on one node of 4 x NVIDIA H100-SXM5 Tensor Core-GPUs (94 GB HBM2e). Table 10 details the hyperparameters used for finetuning.
Hyperparameter falcon-10B llama-8B mistral-7B qwen-3B smollm-1.7B batch_size 4 8 8 8 8 gradient_accumulation 16 8 8 8 8 learning_rate 2.0e-05 1.0e-05 5.0e-06 5.0e-06 5.0e-06 lr_scheduler_type cosine num_train_epochs 2 attn_implementation flash_attention_2 warmup_ratio 0.03 max_seq_length 2048 weight_decay 0.1 neftune_noise_alpha 5 use_liger True
Table 10: Hyperparameter details for finetuning.
A.7 Evaluation – details
In the following, we list the benchmarks and the respective evaluation details, such as few-shot settings and other information about formatting and response extraction.
Benchmark Category ($\mathcal{L}$) Evaluation details IFeval (Zhou et al., 2023b) Generation 0-shot; prompt level strict acc GPQA (Rein et al., 2024) Multiple 0-shot BBH (Suzgun et al., 2022) Reasoning 3-shot; multiple-choice MuSR (Sprague et al., 2024) Reasoning 0-shot; multiple-choice Math (Hendrycks et al., 2021c) Math 4-shot ARC-C (Clark et al., 2018) Multiple 0-shot; multiple-choice GSM8k (Cobbe et al., 2021) Math 5-shot Hellaswag (Zellers et al., 2019) Multiple 0-shot; multiple-choice MMLU (Hendrycks et al., 2021b) Factual QA 0-shot; multiple-choice TruthfulQA (Lin et al., 2021) Factual QA 0-shot; multiple-choice Winogrande (Keisuke et al., 2019) Multiple 0-shot; multiple-choice MBPP (Austin et al., 2021) Coding 3-shot; pass@1 OpenBookQA (Mihaylov et al., 2018) Extraction 0-shot; multiple-choice
Table 11: List of used benchmarks with associated category and evaluation details.
A.8 Results – details
A.8.1 Results – scaling by benchmark
We present here the results of our scaling experiments, without aggregating them across evaluation benchmarks.
(a) ARC-challenge (Clark et al., 2018)
(b) GSM8k (Cobbe et al., 2021)
(c) HellaSwag (Zellers et al., 2019)
(d) BBH (Suzgun et al., 2022)
(e) GPQA (Rein et al., 2024)
(f) IFeval (Zhou et al., 2023b)
(g) Math (Hendrycks et al., 2021c)
(h) MuSR (Sprague et al., 2024)
(i) MBPP (Austin et al., 2021)
(j) MMLU (Hendrycks et al., 2021b)
(k) TruthfulQA (Lin et al., 2021)
(l) Winogrande (Keisuke et al., 2019)
Figure 11: Results scaling across all benchmark datasets
Given the following categories:
- Math (math questions and math reasoning problems)
- Coding (programming tasks or coding questions)
- Generation (creative generation tasks with constraints, including roleplaying)
- Reasoning (logical deductive reasoning tasks that are neither math nor coding)
- Brainstorming (information-seeking or recommendation questions requiring explanation, or classification tasks)
- Factual QA (simple factual questions, without any context)
- Extraction (extraction tasks, including QA, from a given textual passage)
What is the category of the following task? Please respond only in JSON format (e.g., {{"answer": "Generation"}})
### Task ###
{input}
Figure 12: Zero-shot prompt for instruction categorization.
### Task ###
1. Given the following user’s PROMPT and system’s RESPONSE, please review the code snippet in the RESPONSE.
2. Focusing on functional correctness, give the final verdict: ’correct’ vs ’incorrect’.
3. Extract the original code snippet, write "no code" if there’s no code snippet.
4. If the original code is correct, simply write "no revision", otherwise propose a code revision to improve the code.
5. Provide your answer in JSON format, with "review", "final_verdict", "code_original" and "code_revision" as keys.
### User’s PROMPT ###
{instruction}
### System’s RESPONSE ###
{output}
Figure 13: Zero-shot prompt for code review and code revision.
### Task ###
1. Given a USER’s prompt, decide whether the constraints from the list below are expressed in the USER’s prompt (yes/no).
2. Provide the expressed constraints in JSON format with the expressed constraint as the key
   and the constraint type as the value if the respective constraint is expressed.
### List of Constraints ###
- letter_case, e.g., lowercase, all capitals
- placeholder_and_postscript
- repeat_prompt, e.g., repeat the request
- output_combination, e.g., multiple responses, separate the response
- choose_output, e.g., choose answer from given options
- output_format, e.g., json format, markdown format, bulleted list, formatted title, highlighted sections
- keyword_included, e.g., included words
- keyword_avoided, e.g., avoided words
- keyword_frequency, e.g., five hashtags, ’but’ two times, letter ’r’ at least 3 times
- language, e.g., english, two languages
- length, e.g., number of words, number of sentences, number of paragraphs
- punctuation, e.g., no commas, quotation
- start_and_ending, e.g., start with ’Hello’, end with ’Thank you!’
- writing_style (e.g., shakespeare, easy-to-read, 5-year-old, persuasive)
- writing_type (e.g., letter, email, proposal, poem)
- topic (e.g., love)
### Examples ###
{few-shot_examples}
### Question ###
USER: {instruction}
Figure 14: Few-shot prompt for constraint identification.
### Task ###
Answer the following questions. Provide the answer in JSON format with the question number as the key and the answer as
the value (true/false).
Questions:
{questions}
ASSISTANT:
{output}
Figure 15: Zero-shot prompt for constraint verification.
### Task ###
Given the USER’s prompt and ASSISTANT’s response below, analyze whether the response addresses USER’s intents properly,
while respecting any constraints expressed in the prompt. Based on these judgments, provide the final score of response
quality in the range of 1 to 10, in JSON format (’score’ as key and quality score as value).
USER: {instruction}
ASSISTANT:
{output}
Figure 16: Zero-shot prompt for response evaluation.