### v3

# Laser: Stratified Selective Sampling for Instruction Tuning with Dedicated Scoring Strategy

Abstract
Recent work shows that post-training datasets for LLMs can be substantially downsampled without noticeably deteriorating performance. However, data selection often incurs high computational costs or is limited to narrow domains. In this paper, we demonstrate that data selection can be both—efficient and universal—by using a multi-step pipeline in which we efficiently bin data points into groups, estimate quality using specialized models, and score difficulty with a robust, lightweight method. Task-based categorization allows us to control the composition of our final data—crucial for finetuning multi-purpose models. To guarantee diversity, we improve upon previous work using embedding models and a clustering algorithm. This integrated strategy enables high-performance fine-tuning with minimal overhead.
Laser: Stratified Selective Sampling for Instruction Tuning
with Dedicated Scoring Strategy
Paramita Mirza1, Lucas Weber2, Fabian Küch2 1ScaDS.AI & TU Dresden, Germany, 2Fraunhofer IIS, Germany paramita.mirza@tu-dresden.de    {lucas.weber, fabian.kuech}@iis.fraunhofer.de
footnotetext: Accepted at Findings of the Association for Computational Linguistics: EMNLP 2025.**footnotetext: These authors contributed equally to this work.
1 Introduction
Large language models (LLMs) can perform a wide range of text-based tasks through chat interfaces. Their generalist abilities stem from an extensive post-training phase, during which they are optimized to generate useful responses to user queries (Sanh et al., 2022; Mishra et al., 2022; Wei et al., 2022; Ouyang et al., 2022). The choice of training data in this phase has a major impact on model performance (e.g. Zhou et al., 2023a; Xia et al., 2024; Liu et al., 2024b; Chen et al., 2024; Ge et al., 2024).
Prior work identifies three key properties of effective instruction-tuning (IT) data (Zhou et al., 2023a): i) difficulty/relevance, reflecting how much a query contributes to learning (Li et al., 2024b; Liu et al., 2024a, b; Zhao et al., 2024b); ii) quality, the usefulness and accuracy of responses (Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b); and iii) diversity, the scope of within-domain variability and cross-domain coverage of the data (Ge et al., 2024; Lu et al., 2024). Although many approaches address one or more of these aspects, they often face limitations that make them difficult to apply in practice.
Figure 1: Overview of our three-stage pipeline, Laser. Instructions are first classified into seven types, then routed and scored for difficulty ($f$) and quality ($q$). Their combined scores ($p$) are ranked, and top examples are sampled both within embedding clusters and overall, while preserving fixed category proportions.
Most existing methods cover only a subset of criteria (e.g., diversity in Chen et al. 2024; quality and diversity in Ge et al. 2024). To our knowledge, Deita (Liu et al., 2024b) is the only complete pipeline that integrates all three simultaneously. Beyond completeness, prior work can be further refined in terms of generality and robustness. Many heuristic-based methods cannot be applied to different tasks (Muennighoff et al., 2025) or domains (Zhou et al., 2023a). Broader approaches—that do not limit themselves explicitly to a single domain—are usually evaluated only on narrow domains or task-types (e.g., open-ended language generation), fine-tuned on a small set of models (typically from the Llama or Mistral family with 3B–8B parameters), or limited to a single source dataset. Outside these settings, they often prove brittle, failing to consistently outperform random sampling or simple heuristics such as a longest baseline (Diddee and Ippolito, 2025; Zhao et al., 2024a). Ultimately, many methods are expensive as they rely on large or closed-weight LLM judges (Chen et al., 2024; Lu et al., 2024).
We address these limitations with Laser (Label-Aware Scoring and clustERing), a robust, complete, and efficient data selection pipeline. [^1] Code, model access, and replication details will be released upon publication at https://github.com/Fraunhofer-IIS/LASER
Our key contributions are the following:
  1. (i)
    A novel, complete and robust data sampling strategy for instruction-tuning that integrates efficient scoring of difficulty and quality via routing while maintaining both inter- and intra-class diversity through stratified sampling and clustering;
  2. (ii)
    A new and general approach to estimate the difficulty of instructions;
  3. (iii)
    Task-specific quality scoring strategies, including custom-designed scorers, particularly for constrained generation and coding tasks focusing on responses’ correctness; and
  4. (iv)
    A large-scale evaluation spanning diverse model families and sizes, on a wide range of benchmarks covering various task types, with comparisons to strong baselines.
2 Related work
IT data selection.
Prior work on IT data selection generally falls into two categories based on how they assess sample difficulty, quality, and diversity: external scoring and model-inherent criteria. External scoring methods rely on: (i) handcrafted features such as coherence, grammaticality, naturalness and understandability (Cao et al., 2024); (ii) heuristics based on length or formatting (Zhao et al., 2024a; Muennighoff et al., 2025); and (iii) scores derived from LLMs of varying scales (LLM-scorers). Notable approaches leveraging LLM-scorers are: AlpaGasus (Chen et al., 2024), which prompts ChatGPT to score each data point; InsTag (Lu et al., 2024), which uses ChatGPT to generate open-ended tags for deriving complexity/diversity; Deita (Liu et al., 2024b), which fine-tunes LLaMA-1-13B on ChatGPT-annotated complexity/quality labels as Deita-scorers; and CaR (Ge et al., 2024), which trains a 550M reward model to rank instruction-response pairs by quality.
On the other hand, model-inherent criteria form another group of approaches that rely on signals from the target model itself, including: LESS (Xia et al., 2024), which estimates data influence using gradient information; SelectIT (Liu et al., 2024a), which measures uncertainty via token probability, prompt variation and multiple models’ assessments; SHED (He et al., 2024), which clusters data and estimates impact using Shapley values; instruction-following difficulty (IFD; Li et al., 2024b), measuring discrepancies between the model’s intrinsic generation capability and its desired response; and the approach by Li et al. (2024c), which evaluates sample utility based on how much it reduces loss when used as an in-context example.
Our approach embraces external and model-inherent perspectives simultaneously: we use domain-specific LLM-scorers to assess quality, and leverage various models’ performances across benchmarks to estimate difficulty—that is, how likely an average model is to fail on a given instruction. For diversity, our approach closely follows CaR (Ge et al., 2024), which clusters data points and samples one representative from each cluster.
IT data categorisation.
Previous work has proposed various task types (e.g., open QA, brainstorming, creative writing), to guide IT data collection from human annotators (Ouyang et al., 2022; Conover et al., 2023). Grattafiori et al. (2024) finetuned Llama 3 8B for coarse-grained (e.g., mathematical reasoning) and fine-grained (e.g., geometry and trigonometry) topic classification to help filter low-quality samples, though they provide little detail on the methodology or intended downstream use. Other efforts impose tags or domain taxonomy on IT data to ensure diversity (Lu et al., 2024; Muennighoff et al., 2025). Dong et al. (2024) study how data composition across tasks affects model performance, but assume that ShareGPT only contains general alignment tasks, while our analysis shows that nearly 30% of it actually consists of coding tasks. To our knowledge, no prior work leverages IT data categorisation to apply task-specific scoring strategies during selection.
Utilities of IT data selection methods.
Several papers criticized the effectiveness and cost-efficiency of existing IT data selection strategies. Zhao et al. (2024a) show that simply selecting the longest responses can outperform more complex methods while being significantly cheaper and easier to implement. Similarly, Diddee and Ippolito (2025) find that many sophisticated methods barely outperform random sampling under realistic conditions, and emphasize the cost-performance trade-off. However, most of these comparisons are limited to a single-source sampling setup, whereas a more practical scenario involves selecting from a pool of IT data sources. Additionally, prior evaluations focus on LLaMA models with a few exceptions using Mistral Liu et al. (2025), leaving the generalizability of selection strategies across model families and sizes largely unexplored.
3 Methods
Our goal is to determine a subset of an arbitrarily large instruction source dataset that is effective when finetuning a given pre-trained base model. We define an effective dataset as one which achieves high performance on a large and general set of evaluation tasks, while requiring relatively few model parameter updates. Formally, let $\mathcal{D}=\{d_{i}\}_{i=1}^{N}$ with $d_{i}=(x_{i},y_{i})$ be the full source dataset of size $N$, where $x_{i}$ represents an input sequence (i.e. an instruction) and $y_{i}$ represents the corresponding output (i.e. the desirable model response). We wish to select a subset $\mathcal{D}^{\prime}\subset\mathcal{D}$ of size $m$ (with $m\ll N$) that is most effective for instruction tuning.
Previous research has shown that i) instruction difficulty (see e.g. Li et al., 2024b; Liu et al., 2024a, b; Zhao et al., 2024b) ii) response quality (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b) and iii) diversity/composition (see e.g. Ge et al., 2024; Lu et al., 2024) are crucial for effective data selection for LLM finetuning. We address these three aspects in the three-step Laser pipeline as illustrated in Figure 1:
  1. 1.
    Classification. We train a lightweight classifier $\pi_{c}$ to categorise all inputs $x_{i}$ in $\mathcal{D}$ into one out of seven categories, which we denote as $l\in\mathcal{L}$.
  2. 2.
    Scoring. Each input–output pair $(x_{i},y_{i})$ is scored using a general-purpose difficulty scorer (yielding $f_{i}$) and a category-specific quality scorer (yielding $q_{i}$); the results are combined into an overall preference score $p_{i}$.
  3. 3.
    Clustering + Ranking. Ultimately, we select the samples with the highest $p_{i}$ while using a clustering approach to maintain diversity and minimise redundancies in $\mathcal{D}^{\prime}$.
We detail each step in the remainder of this section.
3.1 Instruction Classification
Inspired by the use-case categories defined in Ouyang et al. (2022), we establish the following task categories $\mathcal{L}$ for samples in instruction tuning datasets: i) Math, from simple calculation to problems requiring multi-step reasoning; ii) Coding, code generation tasks or programming-related question answering; iii) Generation, textual generation tasks including roleplaying, summarizing and rewriting passages; iv) Reasoning, questions requiring deductive/logical reasoning; v) Brainstorming, information-seeking and recommendation questions that require inductive reasoning, including classification tasks; vi) Factual QA, factual questions with simple facts as answers; and vii) Extraction, tasks requiring structured/answer extraction from textual contexts. We let two human annotators classify 80 samples from the popular MT-Bench dataset (Zheng et al., 2023) and achieve high inter-annotator agreement (Cohen’s Kappa = 0.8635), demonstrating the discriminability of our categories. We compare two different approaches to build a classification model: i) LLM annotator and ii) SetFit classifier.
With the LLM-annotator approach (Wei et al., 2022), we prompt instruction categorization by listing categories with brief explanations, followed by “What is the category of the following task?” (see Figure 19, Appendix A.1). Meanwhile, SetFit (Tunstall et al., 2022) is a few-shot learning method that tunes Sentence Transformers (Reimers and Gurevych, 2019) on labelled input pairs in a contrastive, Siamese manner. We manually identify approximately 250 samples strongly associated with each category (see Table 12, Appendix A.1) to train the SetFit classifier; hyperparameters are detailed in Appendix A.1.
For evaluation, we use the manually annotated MT-Bench dataset as the test set, assessing accuracy, macro F1-score, and Cohen’s Kappa agreement with human judgment (see Table 1). While zero-shot prompting with GPT-4o performs best, we choose the SetFit classifier for our pipeline due to its comparable performance and higher efficiency than larger LLMs like GPT-4o and Falcon3-10B-Instruct (Team, 2024b).
Approach LLM annotator/embedding model Acc F1 Cohen’s Kappa Zero-shot GPT-4o 0.88 0.86 0.85 prompting tiiuae/Falcon3-10B-Instruct 0.85 0.84 0.82 meta-llama/Llama-3.1-8B-Instruct 0.76 0.65 0.71 SetFit NovaSearch/stella_en_400M_v5 $\dagger$ 0.85 0.81 0.82 classifier Lajavaness/bilingual-embedding-large 0.82 0.78 0.79 NovaSearch/stella_en_1.5B_v5 0.66 0.60 0.60
Table 1: Evaluation results on instruction categorization. $\dagger$ denotes the chosen approach and model for our data selection pipeline.
3.2 Scoring
3.2.1 Difficulty scorer
Previous research suggests that difficulty (sometimes referred to as complexity) matters for data selection (see e.g. Liu et al., 2024b; Cao et al., 2024; Zhao et al., 2024b; Muennighoff et al., 2025), with more challenging data generally resulting in better model performance. However, existing difficulty metrics either lack generality across domains (e.g. length of reasoning trace in response in Muennighoff et al., 2025, or the Deita complexity scorer222hkust-nlp/deita-complexity-scorer as shown in Figure 10(b)) or are strongly influenced by spurious features (e.g. the widely used Deita-complexity is strongly biased towards long sequences; see Figure 11; Liu et al., 2024b).
Our goal is to train a general and robust difficulty scorer that predicts how likely it is for an average model to solve a data point incorrectly, independent of its category $l$. We denote this difficulty score as $f$.
To source the training set $\mathcal{D}_{\text{diff}}$ for such a scorer, we collect 20k instruction-response pairs, approximately equally distributed across categories $l_{i}\in\mathcal{L}$. We list the data sources in Table 13 and show the category proportions in Figure 8. Then, we evaluate every item with a heterogeneous pool of 18 instruction‑tuned LLMs (see Table 4). We continue by applying multiple preprocessing steps to the model scores: First, we normalise the item scores to the interval $[0,1]$ and remove items with a score of 0 across all models, as they are likely to contain noise or annotation errors and will not yield any valuable learning signal. We then go on to convert the absolute scores into relative deviations with respect to the model’s mean performance on the corresponding dataset, by subtracting the average from the absolute score on each item (illustrated in Figure 9). This step has the effect of weighing wrong responses of strong models more than wrong responses of weaker models (and vice versa). Further, it mitigates potential skews in the model performance distribution (e.g. if many models perform weakly on a specific dataset). Ultimately, we obtain difficulty targets $f$ by averaging over the model pool for every item.
We fine‑tune a Qwen‑3‑8B backbone (Team, 2025) with a single‑layer regression head to minimise the mean‑squared error on this dataset (training details can be found in Appendix A.2.2).
Difficulty scorer evaluation.
We evaluate the effectiveness and the validity of the difficulty scoring. For effectiveness, we sample 25k data points from $\mathcal{D}$ (as it is described in Section 4.1) either at random or following the difficulty scores. We then finetune Mistral-7B-v0.3 (Jiang et al., 2023) on the resulting datasets. Appendix A.2.3 shows how the difficulty scorer helps improve performance on benchmarks independently of their domain. For validity (i.e. whether we are really measuring some notion of difficulty), we score the CodeForces section of DeepMind’s code-contests dataset (Li et al., 2022) and find high correlations with its human-annotated difficulty scores (details in Appendix A.2.4).
3.2.2 Quality scorer
Prior work shows that sample selection strategies based on response length or quality, as judged by external models, lead to better instruction tuning datasets (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b). However, these strategies underestimate the diverse problem types within instruction tuning data, which may require different evaluation criteria. For example, the quality of responses to math and coding problems is heavily dependent on solution correctness, while constrained generation requires evaluation of adhered constraints. In this work, we designate a dedicated quality scorer for each category $l\in\mathcal{L}$ defined in Section 3.1.
Deita quality scorer.
For Reasoning, Factual QA and Extraction samples, we employ the Deita quality scorer333hkust-nlp/deita-quality-scorer (Liu et al., 2024b), a finetuned LLaMA-13B for quality assessment, which yields the $q_{\mathrm{deita}}$ score.
Process reward model.
For Math data, we assess the soundness of mathematical reasoning using a specialised process reward model (Qwen2.5-Math-PRM-7B; Zhang et al., 2025). Process reward models (PRMs; Lightman et al., 2024; Uesato et al., 2022) are trained to verify steps in reasoning traces as they are common in mathematical reasoning. We find double linebreaks and—if no double linebreaks present—single linebreaks as a delimiter to be a good heuristic to separate reasoning steps. As a reasoning trace breaks with a single erroneous step, we aggregate scores by taking the minimal score out of all steps within each trace, as the $q_{\mathrm{math}}$ score.
Code quality scorer.
We design a quality scoring framework for Coding samples by drawing inspiration from Wadhwa et al. (2024). For each data point $(x_{i},y_{i})$, we leverage code-oriented LLMs to: (i) assess the functional correctness of the code snippet in $y_{i}$ with respect to the problem $x_{i}$, and (ii) produce a revised version that improves or fixes the original code (see Figure 20, Appendix A.3.2). The resulting score, $q_{\mathrm{code}}$, is based on the normalized Levenshtein similarity between lines of the original ($lo_{0},...,lo_{n}$) and revised ($lr_{0},...,lr_{m}$) code: $nls=(\max(n,m)-\mathrm{lev}(lo,lr))/\max(n,m)$, where $\mathrm{lev}(lo,lr)$ is the line-level Levenshtein distance. If the original code is functionally correct, we set $q_{\mathrm{code}}=nls$; otherwise, $q_{\mathrm{code}}=nls/2$. If no code snippet is present, we assign $q_{\mathrm{code}}=0.5$ if $y_{i}$ is judged correct, and 0.0 otherwise.
To evaluate this scoring method, we use a 1K-sample test set from LiveCodeBench (Jain et al., 2024a), containing coding problems and LLM-generated responses.444livecodebench/code_generation_samples Using Qwen/Qwen2.5-Coder-14B-Instruct as the reviewing model, our framework achieves 70% accuracy and a 0.412 Pearson correlation with binary correctness labels (see Appendix A.3.2 for details).
Instruction-following scorer.
For Generation and Brainstorming data, we implement an if-quality scorer. Influenced by the IFEval benchmark (Zhou et al., 2023b) which defines “verifiable constraints” such as length (“400 or more words”) and keyword (“without using the word sleep”) constraints, we design a response quality scorer based on the fraction of expressed constraints ($\mathcal{C}_{exp}$) adhered by the response ($\mathcal{C}_{true}$). First, we use an LLM annotator to identify $\mathcal{C}_{exp}$, which comprises (span, constraint type) pairs $\{(s_{i},c_{i})\}^{n_{exp}}_{i=1}$, with $s_{i}$ represents the textual span found and $c_{i}$ is the corresponding constraint label. For example, given the prompt “Write a funny blog post with 400 or more words about the benefits of sleeping in a hammock, without using the word sleep.”, $\mathcal{C}_{exp}=$ {(400 or more words, length), (without using the word “sleep”, keyword avoided), (funny blog post, writing type)}. A list of considered constraint types is shown in Figure 21 (Appendix A.3.1). Next, $\mathcal{C}_{exp}$ is passed to a constraint checker module, which performs two steps:
  1. 1.
    Heuristic verification: We verify length, letter case, punctuation and keyword constraints, by adapting the IFEval verification script.
  2. 2.
    LLM-judge verification: We ask an LLM judge to assess constraints that cannot be verified heuristically (e.g., “Does the following text follow the [writing type] constraint of [funny blog post]?”).
This yields $\mathcal{C}_{true}=\{(s_{j},c_{j})\}^{n_{true}}_{j=1}$, with which we compute the quality score as $q_{\mathrm{if}}=n_{true}*(n_{true}/n_{exp})$, giving more incentives to responses adhering to more constraints. If $\mathcal{C}_{exp}$ is empty, we ask an LLM judge to evaluate whether the response i) addresses the user’s intent, while ii) respecting any constraints expressed in the prompt, and to provide a final scorescore (1–10), which we use to compute the score as $q_{\mathrm{if}}=score/10$.
Our analysis with the IFEval benchmark dataset containing sample responses from ten models as our test bed (see Table 6, Appendix A.3.1), shows that Qwen3-14B (Team, 2025) outperformed other medium-sized Instruct-LLMs as both LLM annotator and judge (see Table 7, Appendix A.3.1). It achieved a macro F1-score of 0.86 for identifying expressed constraints, a Pearson correlation coefficient of 0.523 at the instance-level, and 0.995 at the model-level, where it effectively replicated the IFEval model ranking.
Quality scorer evaluation.
Similar to the difficulty scorer, we demonstrate the effectiveness of our task-specific quality scorers for ranking-based data selection, particularly when compared to the general-purpose Deita quality scorer and random sampling. The corresponding results are presented in Appendix A.3.3.
Dataset #samples (#turns) HuggingFaceH4/ifeval-like-data 5K vicgalle/alpaca-gpt4 (Taori et al., 2023) 52K nvidia/OpenMathInstruct-2 (w/o augmented problems) (Toshniwal et al., 2024) 52K ai2-adapt-dev/flan_v2_converted (Longpre et al., 2023) 90K openbmb/UltraInteract_sft (Coding) (Yuan et al., 2024) 115K WizardLMTeam/WizardLM_evol_instruct_V2_196K (Xu et al., 2024) 143K theblackcat102/sharegpt-english 50K (392K) microsoft/orca-agentinstruct-1M-v1 (random subset) (Mitra et al., 2024) 200K (903K) all 707K (1.75M)
Table 2: Source dataset $\mathcal{D}$
Figure 2: Category distribution in IT datasets with varied sampling.
3.3 Sampling
Overall preference scores.
For each $(x_{i},y_{i})\in\mathcal{D}$, we compute the preference score $p_{i}=f_{i}\cdot q_{i}$, where $f_{i}$ and $q_{i}$ are the difficulty and quality scores, respectively. $f_{i}$ and $q_{i}$ are normalized per scorer as they have differing ranges, using min-max scaling, with the 1st and 99th percentiles as the minimum and maximum values across all samples in $\mathcal{D}$. For multi-turn conversations, where each data point $d_{i}$ consists of a sequence of turns $\{(x_{0},y_{0}),...,(x_{T},y_{T})\}$, these scores are averaged across all turns to yield the overall conversation-level scores $f_{i}$ and $q_{i}$.
Clustering + Ranking.
Greedily choosing the highest-scoring samples often leads to redundancy in some domains and under-representation in others. Thus, maintaining diversity in the final dataset $\mathcal{D}^{\prime}$ is essential. Since diversity is a property of the dataset as a whole—not of individual samples—selection should consider the dataset globally (e.g., Ge et al., 2024), rather than relying on iterative, sample-by-sample strategies (e.g., Liu et al., 2024b; Bukharin et al., 2024). Diversity can be promoted top-down by balancing category proportions (see e.g. Grattafiori et al., 2024; Dong et al., 2024), or bottom-up by ensuring sufficient semantic dissimilarity among selected samples (e.g., Liu et al., 2024b; Ge et al., 2024; Lu et al., 2024).
[…]
Given the large value of $m_{l}$ and the high-dimensional nature of the input embeddings, we employ MiniBatchKMeans from scikit-learn with the k-means++ initialization method and a batch size of 2048.
Sampling evaluation.
We evaluate the effectiveness of our stratified sampling and clustering as we did for the difficulty and quality scorer evaluation in Sections 3.2.1 and 3.2.2. In addition, we demonstrate the efficiency and robustness of Laser’s sampling technique, particularly in comparison with Deita (Liu et al., 2024b) and CaR (Ge et al., 2024). The respective results are reported in Appendix A.4.1.
4 Experiments
4.1 Experimental Setup
Datasets.
We use the aggregation of all listed datasets in Table 2 as the source dataset $\mathcal{D}$ in all experiments, if not specified otherwise.
Models.
While in most experiments we finetune Mistral-7B-v0.3, we also demonstrate generalisation across models of varying sizes and families: i) tiiuae/Falcon3-10B-Base, ii) meta-llama/Llama-3.1-8B, iii) mistralai/Mistral-7B-v0.3, iv) Qwen/Qwen2.5-3B and v) HuggingFaceTB/SmolLM2-1.7B.
(a) Score gains per benchmark category
(b) Score gains per model size
Figure 3: Performance gains over the base model for different sampling strategies with 100k samples. We aggregate the results (a) across all benchmarks separated by models; (b) across all models separated by benchmark categories.
Finetuning.
We fine-tune each base model using the SFTTrainer from the Transformer Reinforcement Learning (TRL) library [^6] https://huggingface.co/docs/trl/sft_trainer (von Werra et al., 2020). Model-specific hyperparameters (e.g., learning rate) are selected for each model family and detailed in Appendix A.6. We also employ NEFTune (Jain et al., 2024b), a technique that improves the performance of chat models by injecting noise into embedding vectors during training.
Baselines.
We compare against the following alternative methods for constructing $D^{\prime}\subset D$:
  1. i)
    random—sampling uniformly at random;
  2. ii)
    longest—including samples having the longest responses; [^7] In the case of multi-turn samples, we only consider the last turn’s responses.
  3. iii)
    deita (Liu et al., 2024b)—ranking data points according to complexity and quality scores ($o_{deita}$ and $q_{deita}$, resp.), and iteratively builds $D^{\prime}$ by adding dissimilar samples based on embedding distances.
[…]
Table 3: Evaluation benchmarks with associated category and evaluation details.
Evaluation.
We evaluate the instruction-tuned models on a suite of benchmarks listed in Table 3, leveraging popular evaluation frameworks LM-evaluation harness (Gao et al., 2024) and AlpacaEval [^8] We use openai/gpt-oss-120b as the LLM-judge. (Dubois et al., 2024). We report normalized score gain as the main evaluation metric, computed as the performance improvement over base models, scaled with min-max normalization.
4.2 Results
4.2.1 Main results
For each selection strategy, we construct a subset $\mathcal{D}^{\prime}={d}_{i=1}^{m}\subset\mathcal{D}$ of size $m=100,000$, and subsequently fine‑tune each of the five base models on $\mathcal{D}^{\prime}$. Figure 2 shows the resulting category distributions across sampling strategies, including the baselines. Figure 3(a) reports normalized score gains over the untuned base models for all evaluation benchmarks and models. Laser achieves the highest overall performance, surpassing all baselines and even the models trained on the entire source set ($|\mathcal{D}|>707\text{k}$). We examine the robustness by breaking the aggregated results down by model size (Figure 3(b)). Laser yields the most consistent gains of all tested conditions for all five bases and outperforms the random sampling in all but the smallest model, confirming that the proposed sampling approach transfers well across models.
Figure 4: Average score gains over the base model when ablating Laser with 100k samples: with only $f$, $q$ and $p$ as ranking criteria.
To better understand the contribution of each Laser component, we ablate sampling by ranking with only difficulty scores ($f$), quality scores ($q$), or their combination (pp)—without clustering or stratification—and compare these variants to the full pipeline. As shown in Figure 4, using $f$ or $q$ alone yields clear improvements over random sampling. However, their full potential is realized only when combined within the complete Laser pipeline. Notably, relying solely on $p$ for ranking provides only marginal gains, and even underperforms $f$- or $q$-based sampling individually, highlighting the importance of diversifying the sampling through stratification and clustering in Laser.
Hence, all three properties (difficulty, quality and diversity) are necessary but not sufficient for sampling data that consistently performs well. This is intuitive: highly challenging queries answered incorrectly provide little benefit, while accurate responses to trivial questions yield weak training signals. Another salient result is the high performance of $q$-sampling for Coding and Math. While the soundness of responses (i.e. $q$-scores) is arguably more important in these domains, we also find that $q$- and $f$-scores in our source data $\mathcal{D}$ are generally negatively correlated, with especially high negative correlations in Coding (r=.38r=-.38) and Math (r=.43r=-.43). Shifting emphasis onto the difficulty ($f$-scores and $p$-scores) might therefore reduce the response quality in these domains and explain the performance gaps observed in Figure 4. We find similar relationships between quality and difficulty for the source datasets $\mathcal{D}_{weak}$ and $\mathcal{D}_{strong}$ in the following section.
4.2.2 Additional robustness experiments
We continue to demonstrate the robustness of the full Laser pipeline for different target sample sizes $m$ and variations in the source data distribution $\mathcal{D}$, such as differences in source data quality and skewed data distributions.
Downscaling $m$.
As different users may require instruction data sets of different sizes, a selection pipeline should improve sampling at every target scale. To verify this, we fix the base model to Mistral‑7B‑Base and constructs datasets $\mathcal{D}^{\prime}$ of {1k, 5k, 10k, 25k, 50k, 100k} items using Laser, as well as random and all data as baselines. Figure 5 demonstrates that our method outperforms the alternatives across scales, except when $m=1,000$. Notably, with only $m=100,000$ (14% of the source data), it surpasses the performance of full-data fine-tuning. We provide benchmark-specific plots in Figure 24 (Appendix A.7), reported without normalizing score gain.
Figure 5: Average score gains over the base model, when finetuning Mistral-7B-Base using instruction datasets with varying sample sizes.
Effectiveness shifts in distribution of $\mathcal{D}$.
Next, it is very probable that the average effectiveness of the source data $\mathcal{D}$ changes from scenario to scenario. We therefore test Laser with two shifted source distributions: $\mathcal{D}_{\text{strong}}$ consisting of the full Tülu v3 (Lambert et al., 2024)—a meticulously optimised dataset, which can be considered highly effective—and $\mathcal{D}_{\text{weak}}$ composed of multiple datasets that have previously proved to be less effective for instruction tuning (see Appendix A.5.2 for details).
We sample $m\in\{1\text{k},5\text{k},10\text{k},25\text{k},50\text{k},100\text{k}\}$ data points from both sources $\mathcal{D}_{\text{strong}}$ and $\mathcal{D}_{\text{weak}}$ using Laser and random sampling. The results of finetuning Mistral-7B-v0.3 are shown in Figure 6. Generally, Laser performs robustly on different source datasets, with fewer fluctuations when changing the sampling size and comparably more improvements for the stronger than for the weaker source dataset. Interestingly, the performance gains using Laser over random are more pronounced for $\mathcal{D}_{\text{weak}}$ when sampling smaller target datasets, potentially distilling the few better examples from the weaker source.
Figure 6: Performance on source datasets of different effectiveness ($\mathcal{D}_{\text{weak}}$ and $\mathcal{D}_{\text{strong}}$) for different sample sizes. For comparison, “all” shows performance when using full $\mathcal{D}_{\text{weak}}$ or $\mathcal{D}_{\text{strong}}$ without sampling.
Figure 7: (a) Category proportions of different $\mathcal{D}_{\text{skewed}}$ source distributions; (b) Average benchmark scores, when finetuning Mistral-7B-Base with data sampled from the skewed distribution using various strategies.
Skewed distribution of $\mathcal{D}$.
Ultimately, $\mathcal{D}$ might be very homogeneous (i.e. contains mostly a specific data type). To simulate this, we create five skewed source sets $\mathcal{D}_{\text{skewed}}$ by randomly selecting two categories $l_{1},l_{2}\in\mathcal{L}$ for each $\mathcal{D}_{\text{skewed}}$ and retaining only items with labels $l_{i}\in{l_{1},l_{2}}$, along with a small (4%) residue from all other categories, yielding a strongly biased data distribution. We show the proportions of each category for all five new source datasets in Figure 7(a).
From each $\mathcal{D}_{\text{skewed}}$, we sample 25k items with the same strategies as previously and fine‑tune Mistral‑7B‑Base on each of the resulting $\mathcal{D}_{\text{skewed}}^{\prime}$. We expected diversity-aware sampling strategies like Deita and Laser to mitigate bias by enforcing either semantic spread and explicit quotas. Indeed, the diversity-aware sampling strategies on average outperform the naive methods, with Laser having an edge over Deita, in terms of performance improvements as illustrated in Figure 7(b). Upon closer inspection, we find that only Laser reduces the introduced bias of categories whereas Deita exacerbates it (see Appendix A.5.2 for details). Next to that, we observe a key limitation of Deita’s sampling method in this setup: only adding dissimilar examples leads to quick saturation when source data is homogeneous and, thereby, hitting a ceiling of possible samples that can be sampled with Deita.
5 Discussion and Conclusion
Effective data selection for instruction tuning is increasingly important as we handle the increasing amount of available data of mixed quality. Nevertheless, the results of data selection pipelines have often been brittle, struggling to generalise across training setups (Diddee and Ippolito, 2025; Zhao et al., 2024a).
In this paper, we address this challenge by explicitly covering the whole spectrum of possible instruction domains and tailoring scoring strategies that are apt to judge the utility of samples in those domains. We show the robustness of our approach by testing it across diverse settings (model families, scales, various shifts in the source data distribution) and evaluating it on a wide range of common benchmarks. Our experiments provide further evidence for the necessity of data selection, as our sampling not only significantly reduces the amount of data required for training, but also outperforms setups trained on the full source data. Despite its more elaborate design—including multiple scoring methods—our pipeline remains computationally efficient due to targeted scoring.
While our pipeline outperforms baselines in almost all settings, it shows only little improvements when sampling large portions from weak source data (as shown in Section 4.2.2). This may reflect a potential ceiling effect on possible performance given the weak source material, suggesting that improving the data itself—rather than relying solely on elaborate sampling techniques—might be necessary. In such cases, iterative data refinement seems promising, especially when guided by reliable scorers for assessing sample difficulty and quality. The modular setup of Laser allows for further incremental improvements of different scorers in future work.
Limitations
Difficulty scorer.
A major issue in collecting data for our difficulty scorer is a certain unreliability in the evaluation of model responses. Previous research shows that evaluations oftentimes have weak robustness to e.g. the prompt formatting (Weber et al., 2023a, b; Sclar et al., 2024; Polo et al., 2024b), bias of LLMs-as-a-judge (Panickssery et al., 2024; Stureborg et al., 2024; Wataoka et al., 2024) or errors in post-processing (such as issues in extracting the answer from the model response). For example, GPT4o showed overall weaker performance on some of our evaluated subsets than some small open-source models. Upon closer inspection, we encountered that—while providing the correct answer—GPT4o generally tends not to follow the formatting of the given few-shot examples and rather responds in an open-form manner, resulting in failing the tight response search masks of the used evaluation frameworks. While we try to mitigate this issue as much as possible, we cannot guarantee that difficulty scores exactly reflect a model’s capacity to solve a given data point.
Quality scorer.
We did not develop nor have dedicated quality scorers for samples belonging to Reasoning, Factual QA and Extraction categories. However, we hypothesize that process reward models (PRMs) could be adapted to Reasoning tasks beyond math, such as spatial (e.g., Wu et al., 2024) and deductive reasoning (e.g., Seals and Shalin, 2024), given appropriate training data. Factuality assessment is a long-standing research problem that has become more relevant in the era of LLMs. Wei et al. (2024) introduced SAFE (Search-Augmented Factuality Evaluator), an LLM-agent-based method for automatically assessing long-form factuality in model response. Although significantly cheaper than human annotators (up to $20\times$), SAFE still incurs costs of \$20–\$40 per 100 prompt-response pairs. For Extraction tasks, future work might draw inspiration from recent advances in RAG evaluation (Yu et al., 2025). These directions, however, are beyond the scope of this paper.
Performance with weak source data.
Our results indicate modest performance gains in setups with very weak source data. However, there might be a ceiling for possible performance in such settings, where data selection alone might not be sufficient to achieve good performance. In such cases, iterative data refinement appears to be a promising approach, particularly when supported by reliable scorers for assessing sample difficulty and quality.
Acknowledgments
We thank Rishiraj Saha Roy, Lucas Druart and Christian Kroos for their valuable feedback on earlier versions of this manuscript. This work was supported in part by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the OpenGPT-X project (no. 68GX21007D) and in part by the Free State of Bavaria in the DSgenAI project (Grant Nr.: RMF-SG20-3410-2-18-4). We also gratefully acknowledge the Centre for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] and ScaDS.AI at the Technical University of Dresden for providing its facilities for experimental work.
Appendix A Appendix
A.1 Instruction Categorization – details
We present the zero-shot prompt for classifying instructions in Figure 19. As for training the SetFit classifier, we leverage the training data detailed in Table 12, and we employ the following hyperparameters. Note that training with SetFit consists of two phases behind the scenes: finetuning embeddings and training a differentiable classification head. As a result, some of the training arguments can be tuples, where the two values are used for each of the two phases, respectively.
  • batch_size=(16, 2)
  • num_epochs=(1, 15)
  • end_to_end=True (train the entire model end-to-end during the classifier training phase)
  • body_learning_rate=(2e-5, 1e-5), the second value is the learning rate of the Sentence Transformer body during the classifier training phase
  • head_learning_rate=1e-4
  • max_steps=500
A.2 Difficulty Scorer – details
A.2.1 Generating $\mathcal{D}_{\text{diff}}$
Figure 8 shows the proportions of different instruction-response pairs that we use as the training data for the difficulty scorer. In Table 13 we report the benchmark data sources from which we obtained the data, as well as the type of evaluation metric that we use to evaluate the 18 LLMs from Table 4.
Figure 8: Proportions of different categories in the difficulty scorer training set $\mathcal{D}_{\text{diff}}$ after cleaning.
Model Type Size
SmolLM2-135M-Instruct (Allal et al., 2025) Univ. 135M
SmolLM2-360M-Instruct (Allal et al., 2025) Univ. 360M
Qwen2.5-0.5B-Instruct (Team, 2024a) Univ. 0.5B
Qwen2.5-Math-1.5B-Instruct (Yang et al., 2024) Math 1.5B
Qwen2.5-Coder-1.5B-Instruct (Hui et al., 2024) Code 1.5B
Qwen2.5-1.5B-Instruct (Team, 2024a) Univ. 1.5B
SmolLM2-1.7B-Instruct (Allal et al., 2025) Univ. 1.7B
Qwen2.5-Math-7B-Instruct (Yang et al., 2024) Math 7B
Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) Code 7B
Qwen2.5-7B-Instruct (Team, 2024a) Univ. 7B
Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) Univ. 7B
Qwen2.5-14B-Instruct (Team, 2024a) Univ. 14B
Mistral-Small-24B-Instruct-2501 (Mistral AI, 2025) Univ. 24B
Qwen2.5-32B-Instruct (Team, 2024a) Univ. 32B
Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) Code 32B
Qwen3-32B (Team, 2025) Univ. 32B
gpt-4o-mini-2024-07-18 (Hurst et al., 2024) Univ. unk
gpt-4o-2024-08-06 (Hurst et al., 2024) Univ. unk
Table 4: Models that were evaluated to obtain difficulty scores for our difficulty scorer training set.
During the data collection for the difficulty scorer, we collect training sets from different benchmarks as well as data points from OpenAssistent (Köpf et al., 2023) as samples of open-ended generation. We evaluate these samples using the existing evaluation frameworks LM-evaluation harness (Gao et al., 2024), big-code evaluation harness (Ben Allal et al., 2022) and FastChat (Zheng et al., 2023). For the LLM-as-a-judge evaluation, we use GPT4o as the judge and evaluate only a subset of 4 LLMs from Table 4 that we deem representative in terms of capabilities (gpt-4o-2024-08-06, Mistral-Small-24B-Instruct-2501, Mistral-7B-Instruct-v0.3 and SmolLM2-1.7B-Instruct).
Figure 9 demonstrates the preprocessing step of calculating the deviation from the average as mentioned in Section 3.2.1.
Figure 9: An illustration of calculating the deviation of the average per dataset: Assuming we have a strong and a weak model with the average scores of $\bar{s}_{\text{s}}=0.8$ and $\bar{s}_{\text{w}}=0.15$ respectively. We subtract these averages from the respective item scores (here two examples with $s^{+}=1$ and $s^{-}=0$). We can see how an incorrect response of the strong model produces a much more negative score than an incorrect response from the weak model (compare $s_{\text{s}}^{-}$ and $s_{\text{w}}^{-}$). In other words, when a strong model gets an item wrong, this is much more meaningful for its difficulty compared to the weak model. Similarly, when a weak model correctly solves a datapoint, it is a much more meaningful signal for the data point’s ease than when a strong model is correct.
A.2.2 Difficulty scorer training
We equip regular CausalLLMs with a regression head by pooling the final hidden states and adding a linear projection to a scalar output, then finetune these models on the difficulty scoring task using the hyperparameters detailed in Table 5. We evaluate four different base models as difficulty scorer—Llama-3.1-8B (Grattafiori et al., 2024), Qwen-2.5-7B (Team, 2024a), Qwen3-4B and Qwen3-8B (Team, 2025))—and select Qwen3-8B, which achieves the best performance on in-distribution evaluation data.
Hyperparameter Value
batch_size 1
gradient_accumulation 16
learning_rate 1e-5
lr_scheduler_type linear
num_train_epochs 8
warmup_steps 100
max_seq_length 2048
weight_decay 0.01
neftune_noise_alpha 10
Table 5: Hyperparameter details for training the difficulty scorer
A.2.3 Effectiveness
To evaluate the effectiveness of the difficulty scorer for data selection across benchmarks, we sample 25k data points from the datasets in Table 2, either at random or based on Laser difficulty score ranking. The results in Figure 10(a) show that difficulty-based sampling outperforms or is comparable to random sampling on all benchmarks except GPQA. For comparison, Figure 10(b) reports results using the Deita complexity scorer under the same setup.
Figure 10: (a) Performance comparison of difficulty scorer vs. random sampling; (b) Performance comparison of complexity scorer vs. random sampling
Figure 11: Relation between complexity/difficulty scores and length of the input sequence. The length of the input sequence should, in most cases, be unrelated to its difficulty.
Figure 12: Relation between complexity/difficulty scores and the human code-forces difficulty scores. Difficulty scores show higher correlation to human scores than complexity scores.
[…]
A.2.4 Validity
We further evaluate whether the difficulty scores are akin to the notion of difficulty in humans or whether our scorer fits a spurious feature of our training data.
We first test it against instruction length, a common spurious feature of difficulty/complexity scorers. While difficult instructions are plausibly often longer than easier instructions, there is no causal relationship between the length and difficulty. A scorer that is trained to predict difficulty should therefore not rely on input length as a feature to rely its prediction upon. We show in Figure 11 how our difficulty scorer is weakly related to the spurious feature of instruction length, while the correlation of Deita complexity is much higher.
Besides showing how the difficulty scorer does not rely on instruction length, we also use an external criterion to evaluate whether it measures difficulty in the human sense. For this end, we score the CodeForces section of DeepMind’s code-contests dataset (Li et al., 2022) with our difficulty scorer as well as the Deita complexity scorer and correlate the scores with the given cf_rating. Figure 12 shows how the difficulty scorer is much better at predicting human difficulty scores than the Deita complexity scorer.
A.2.5 Experimentation with IRT-based targets
We further experimented with difficulty targets grounded in psychometric methods called item response theory (IRT; see Lord, 1968; Brzezińska, 2020; Van der Linden, 2018; Cai et al., 2016). IRT provides a framework for estimating latent traits from observed responses and has recently been applied to the evaluation of large language models (see e.g. Lalor et al., 2016; Rodriguez et al., 2021; Vania et al., 2021; Zhuang et al., 2023; Polo et al., 2024a). In our setting, an IRT model estimates item-level parameters from correctness scores across the LLMs from Table 4. We fit such a model separately on each dataset in $\mathcal{D}_{\text{diff}}$ and use the resulting item difficulty parameters $\beta$ as targets for training our difficulty scorer. We find that the IRT-based difficulty scorer yields improved results on the correlations with the Codeforces ratings (the validity criterion; see Figure 13). At the same time, this IRT-based difficulty scorer does not yield better results when used for data selection. We see the use of IRT-based difficulty scorers as an exciting future research direction, as the possibilities to gain more detailed insights into data point features are great (e.g. by using 2PL or 3PL instead of 1PL models or by extending from unidimensional to multidimensional models).
A.3 Quality Scorers – details
A.3.1 Instruction-following Scorer
Prompts used by LLM-annotator/judge for deriving instruction-following scores can be found in Figure 21, A.7 and 23. As our test dataset, we collected responses from 10 LLMs on the IFEval benchmark (see Table 6), available on open-llm-leaderboard’s dataset collection of evaluation details. [^9] https://huggingface.co/open-llm-leaderboard Table 7 presents the evaluation results of various LLMs as the annotator/judge.
Model IFEval score meta-llama/Llama-3.3-70B-Instruct 0.90 Qwen/Qwen2.5-14B-Instruct-1M 0.84 allenai/Llama-3.1-Tulu-3-70B-SFT 0.81 tiiuae/Falcon3-7B-Instruct 0.76 ibm-granite/granite-3.1-8b-instruct 0.72 microsoft/Phi-3-medium-128k-instruct 0.60 abacusai/Smaug-34B-v0.1 0.50 Qwen/Qwen2.5-32B 0.41 google/gemma-1.1-2b-it 0.31 databricks/dolly-v2-7b 0.20
Table 6: Performance of 10 considered LLMs on IFEval.
LLM annotator/judge macro Pearson’s $r$ F1 instance-level model-level Qwen/Qwen2.5-7B-Instruct 0.80 0.503 0.986 meta-llama/Llama-3.1-8B-Instruct 0.83 0.454 0.982 tiiuae/Falcon3-10B-Instruct 0.84 0.515 0.969 Qwen/Qwen3-14B $\dagger$ 0.86 0.523 0.995
Table 7: Evaluation results of various LLMs as annotator/judge on identifying expressed constraints (macro-F1), and Pearson correlations coefficient ($r$) of resulting instruction-following scores with IFEval benchmarks scores at both instance-level and model-level. $\dagger$ denotes the chosen LLM annotator/judge for our instruction-following scorer.
LLM reviewer acc Pearson’s $r$ deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct 0.64 0.278 deepseek-ai/DeepSeek-R1-Distill-Qwen-14B 0.66 0.389 Qwen/Qwen2.5-Coder-14B-Instruct $\dagger$ 0.70 0.412 Qwen/Qwen3-14B (non-coding LLM) 0.70 0.411
Table 8: Evaluation results of various LLMs as code reviewer (acc), and Pearson correlation coefficient ($r$) of resulting code quality scores with LiveCodeBench benchmarks binary correctness labels. $\dagger$ denotes the chosen LLM annotator/judge for our instruction-following scorer.
A.3.2 Code-quality Scorer
The prompt used by LLM-annotator/judge for deriving code-quality scores can be found in Figure 20. Table 8 presents the evaluation results of various LLMs as the code reviewer.
Figure 14: Performance of quality scoring-based sampling on various benchmarks.
A.3.3 Effectiveness
We test the efficacy of Laser task-specific quality scorers against the general-purpose Deita quality scorer, by comparing the performance of Mistral-7B-v0.3 finetuned on 25k data points, either randomly selected from data in Table 2, or based on ranking by considered quality scorers. For each considered benchmark, we selected only samples from the relevant category as defined in Table 3. Figure 14 demonstrates the clear gain of process reward model (PRM) scoring for Math benchmarks, particularly GSM8K. For Coding benchmarks, code-quality scoring outperforms Deita quality scoring on HumanEval, but achieves similar results as random. IF-quality scoring surpasses both Deita quality scoring and random on Generation/Brainstorming benchmarks. For other categories (Reasoning, Factual QA and Extraction), Deita quality scoring provides noticeable benefits only on TruthfulQA, with minimal impact on other benchmarks.
Figure 15: Source dataset proportions of different sampling strategies for 100k samples.
A.4 Sampling – details
A.4.1 Effectiveness
In Section 3.3, we introduce a sampling technique that aims to improve the diversity of the target dataset $\mathcal{D^{\prime}}$. The sampling technique consists of two components: firstly, we stratify the sampling by the categories that we assigned using the instruction classifier $\pi_{c}$ (according to the proportions shown in Figure 2 [laser]) and, secondly, we cluster the datapoints in embedding space and restrict the sampling to one datapoint per cluster. We sample here once 25k data points with the first technique (which we will call random+) and once 25k data points with both methods (correspondingly, random++) from $\mathcal{D}$. The random in the method name indicates that we select samples otherwise randomly (hence, not using any scoring during sampling). We then train a Mistral-7B-v0.3-base model on the resulting datasets and compare the results with entirely random sampling in Figure 16.
A.4.2 Efficiency
In Table 9, we compare the sampling runtime of Laser with Deita, which iteratively builds $\mathcal{D^{\prime}}$ by adding dissimilar samples based on embedding distance, as well as CaR, which runs clustering on the entire dataset $\mathcal{D}$. While Deita is highly efficient given small $m$ (e.g., 1k), its runtime grows rapidly with larger $m$. Meanwhile, CaR suffers from out-of-memory issues at large $m$. In contrast, Laser remains stable across different $m$ and is generally more efficient than CaR.
$m$ Deita CaR Laser
1k 92.3 s 5515.7 s 1505.1 s
10k 970.2 s 5750.9 s 1555.9 s
100k 41055.3 s OOM 2818.1 s
Table 9: Runtime comparison for different $m$.
Figure 16: Effect of different sampling strategies on training outcomes. While random selects $m$ = 25k arbitrary samples, random+ samples by stratifying the categories in $D^{\prime}$ according to the proportions shown in Figure 2 (laser). Ultimately, random++ also includes clustering (as described in Section 3.3, while randomly selecting a datapoint from each cluster)
A.5 Datasets – details
A.5.1 Datasets main experiments
Figure 18 shows SetFit classification results for IT datasets from which we composed $\mathcal{D}$. On the other hand, Figure 15 shows which of the source datasets are sampled when we use various sampling strategies. Interestingly, while Laser maintains relative diversity across different source datasets, Deita is completely dominated by only two datasets: microsoft/orca-agentinstruct-1M-v1 (200k sampled) and WizardLMTeam/WizardLM_evol_instruct_V2_196K.
A.5.2 Datasets robustness experiments
Dataset details $\mathcal{D}_{\text{strong}}$ and $\mathcal{D}_{\text{weak}}$
In Section 4.2.2, we introduce $\mathcal{D}_{\text{strong}}$ and $\mathcal{D}_{\text{weak}}$. While $\mathcal{D}_{\text{strong}}$ simply consists of the Tülu v3 dataset, $\mathcal{D}_{\text{weak}}$ is a aggregation of different datasets. The single components exhibited either in our own pilot experiments or in prior related work comparably weak performance to other datasets. This weaker performance also shows, when we compare the finetuning using full $\mathcal{D}_{\text{weak}}$ with the finetuning using full $\mathcal{D}_{\text{strong}}$ in Figure 6, with the latter significantly outperforming the former.
Dataset #samples (#turns) ConiferLM/Conifer (Sun et al., 2024) 13k databricks/databricks-dolly-15k (Conover et al., 2023) 15k akoksal/LongForm (Köksal et al., 2023) 23k alpaca (Taori et al., 2023) 52K vicgalle/alpaca-gpt4 52K ai2-adapt-dev/flan_v2_converted (Sanh et al., 2022) 90K nvidia/Daring-Anteater (Wang et al., 2024) 100k AI-MO/NuminaMath-CoT101010random subset (Li et al., 2024a) 250k all 595K
Table 10: Composition of $\mathcal{D}_{\text{weak}}$
Figure 17: Category distribution for datasets sampled from the 5 bias source datasets $\mathcal{D}_{\text{skewed}}$ using different sampling strategies.
Results of experiments with $\mathcal{D}_{\text{skewed}}$
In Section 4.2.2, we experiment with skewed data distributions. In Figure 17, we show the resulting category distributions when sampling with different sampling approaches from the 5 skewed source distributions. As we can see, only Laser reduces the introduced bias in the skewed source data, while Deita is not effective in reducing the bias, but in many cases exacerbates it.
A.6 Finetuning – details
We finetune all models in our experiments on one node of 4 x NVIDIA H100-SXM5 Tensor Core-GPUs (94 GB HBM2e). Table 11 details the hyperparameters used for finetuning.
Hyperparameter falcon-10B llama-8B mistral-7B qwen-3B smollm-1.7B
batch_size 4 8 8 8 32
gradient_accumulation 16 8 16 8 16
learning_rate 2.0e-05 5.0e-06 5.0e-06 2.0e-05 1.0e-04
num_train_epochs 2 2 2 2 3
weight_decay 0.1 0.01 0.01 0.01 0.01
warmup_ratio 0.03 0.03 0.1 0.03 0.03
lr_scheduler_type cosine
attn_implementation flash_attention_2
max_seq_length 2048
neftune_noise_alpha 5
use_liger True
Table 11: Hyperparameter details for finetuning.
A.7 Results – details
Figure 24 shows the scaling experiment results across benchmarks, reported without normalizing score gain.
Figure 18: Category distribution in IT datasets used in the experiments.
Category Training data (subset) #Samples Math nvidia/OpenMathInstruct-2, AI-MO/NuminaMath-CoT 250 Coding openbmb/UltraInteract_sft (Coding), microsoft/orca-agentinstruct-1M-v1 (code), 253 HuggingFaceH4/no_robots (Coding), lissadesu/codeqa_v3 Generation HuggingFaceH4/no_robots (Generation, Rewrite, Summarize), HuggingFaceH4/ifeval-like-data, 250 declare-lab/InstructEvalImpact (Creative, Professional), iamketan25/roleplay-instructions-dataset Extraction HuggingFaceH4/no_robots (Closed QA, Extract) 250 Factual QA HuggingFaceH4/no_robots (Open QA), basicv8vc/SimpleQA 255 Brainstorming declare-lab/InstructEvalImpact (Informative, Argumentative), HuggingFaceH4/no_robots (Brainstorming, Classify), 250 matt-seb-ho/WikiWhy Reasoning renma/ProofWriter, hitachi-nlp/ruletaker, lucasmccabe/logiqa, lucasmccabe/logiqa, tasksource/strategy-qa 250
Table 12: Training data overview for SetFit classifier.
Dataset Subset Split $n$ samples Eval. metric
GSM8K (Cobbe et al., 2021) Default train 1192 exact match
Math (Hendrycks et al., 2021c) Algebra train 322 exact match
Counting & Probability train 264 exact match
Geometry train 220 exact match
Intermediate Algebra train 233 exact match
Number Theory train 314 exact match
Prealgebra train 348 exact match
Precalculus train 238 exact match
IFEval-like Default - 1990 instance level loose acc
MBPP (Austin et al., 2021) Default train & val 448 pass@1
OpenBookQA (Mihaylov et al., 2018) Default train 299 acc
ARC (Clark et al., 2018) Challenge train 464 acc
bAbI (Dodge et al., 2016) Default train 224 exact match
CommonsenseQA (Talmor et al., 2019) Default train 248 acc
CoQA (Reddy et al., 2019) Default train 380 F1
DROP (Dua et al., 2019) Default train 383 F1
FLD (Morishita et al., 2023) Default train 739 exact match
Logical Formula Default train 750 exact match
HeadQA (Vilares and Gómez-Rodríguez, 2019) English train 689 acc
Spanish train 703 acc
JSONSchemaBench (Geng et al., 2025) Easy train 200 schema compliance & json validity
Medium train 198 schema compliance & json validity
Hard train 179 schema compliance & json validity
LogiQA (Liu et al., 2020) LogiEval train 490 exact match
LogiQA2 train 416 acc
MLQA (Lewis et al., 2020) 49 lang. combinations val 715 F1
TriviaQA (Joshi et al., 2017) Default train 1389 exact match
OpenAssistent (Köpf et al., 2023) Default train 5318 LLM-as-a-judge
APPS (Hendrycks et al., 2021a) introductory train 422 pass@1
interview train 303 pass@1
competition train 289 pass@1
CONALA (Yin et al., 2018) Default train 97 Bleu
Table 13: Datasets used in difficulty scorer training.
Given the following categories:
- Math (math questions and math reasoning problems)
- Coding (programming tasks or coding questions)
- Generation (creative generation tasks with constraints, including roleplaying)
- Reasoning (logical deductive reasoning tasks that are neither math nor coding)
- Brainstorming (information-seeking or recommendation questions requiring explanation, or classification tasks)
- Factual QA (simple factual questions, without any context)
- Extraction (extraction tasks, including QA, from a given textual passage)
What is the category of the following task? Please respond only in JSON format (e.g., "answer": "Generation")
### Task ###
{input}
Figure 19: Zero-shot prompt for instruction categorization.
### Task ###
1. Given the following user’s PROMPT and system’s RESPONSE, please review the code snippet in the RESPONSE.
2. Focusing on functional correctness, give the final verdict: ’correct’ vs ’incorrect’.
3. Extract the original code snippet, write "no code" if there’s no code snippet.
4. If the original code is correct, simply write "no revision", otherwise propose a code revision to improve the code.
5. Provide your answer in JSON format, with r̈eview", "final_verdict", "code_original" and "code_revision" as keys.
### User’s PROMPT ###
{instruction}
### System’s RESPONSE ###
{output}
Figure 20: Zero-shot prompt for code review and code revision.
### Task ###
1. Given a USER’s prompt, decide whether the constraints from the list below are expressed in the USER’s prompt (yes/no).
2. Provide the expressed constraints in JSON format with the expressed constraint as the key
and the constraint type as the value if the respective constraint is expressed.
### List of Constraints ###
- letter_case, e.g., lowercase, all capitals
- placeholder_and_postscript
- repeat_prompt, e.g., repeat the request
- output_combination, e.g., multiple responses, separate the response
- choose_output, e.g., choose answer from given options
- output_format, e.g., json format, markdown format, bulleted list, formatted title, highlighted sections
- keyword_included, e.g., included words
- keyword_avoided, e.g., avoided words
- keyword_frequency, e.g., five hashtags, ’but’ two times, letter ’r’ at least 3 times
- language, e.g., english, two languages
- length, e.g., number of words, number of sentences, number of paragraphs
- punctuation, e.g., no commas, quotation
- start_and_ending, e.g., start with ’Hello’, end with ’Thank you!’
- writing_style (e.g., shakespeare, easy-to-read, 5-year-old, persuasive)
- writing_type (e.g., letter, email, proposal, poem)
- topic (e.g., love)
### Examples ###
{few-shot_examples}
### Question ###
USER: {instruction}
Figure 21: Few-shot prompt for constraint identification.
### Task ###
Answer the following questions. Provide the answer in JSON format
with the question number as the key and the answer as the value
(true/false).
Questions:
{questions}
ASSISTANT:
{output}
Figure 22: Zero-shot prompt for constraint verification.
### Task ###
Given the USER’s prompt and ASSISTANT’s response below,
analyze whether the response addresses USER’s intents properly,
while respecting any constraints expressed in the prompt.
Based on these judgments, provide the final score of response
quality in the range of 1 to 10, in JSON format (’score’ as key
and quality score as value).
USER: instruction
ASSISTANT:
{output}
Figure 23: Zero-shot prompt for response evaluation.
(a) AGIEval (Zhong et al., 2024)
(b) AlpacaEval (Dubois et al., 2024)
(c) ARC-Challenge (Clark et al., 2018)
(d) BBH (Suzgun et al., 2023)
(e) GPQA (Rein et al., 2024)
(f) HellaSwag (Zellers et al., 2019)
[…]
(i) Math (Hendrycks et al., 2021c)
(j) HumanEval (Chen et al., 2021)
(k) MBPP (Austin et al., 2021)
(l) IFeval (Zhou et al., 2023b)
(m) MMLU (Hendrycks et al., 2021b)
(n) MuSR (Sprague et al., 2024)
(o) OpenBookQA (Mihaylov et al., 2018)
(p) TruthfulQA (Lin et al., 2022)
(q) Winogrande (Sakaguchi et al., 2020)
Figure 24: Results scaling across all benchmark datasets